A Complete Bibliography of Publications in the ICES Journal of Marine Science (1970–1979)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

21 August 2024 Version 1.04

Title word cross-reference

2 [Pet77].

0-group [Loc73, LdBD75]. 000 [RG70].

1972 [Lee74]. 1973 [Ram75]. 1975 [Tie76].

40 [Thr70b].

Âge [Du 77].

abdominal [KSL74]. absorption [MS76]. Abundance [Dow72, Pos72, dCR71, Dow77a, Gra78]. Acartia [TM78]. Acoustic [Wes72, AH78, CH76, Cus78, DVB77]. Activity [Saw72, SdG71]. acutifrons [Gue70]. adsorption [TB75]. Adult [GW71, Ile71b, GS74]. aeglefinus [CL78, HH71b, JH72, Jon74, Jon78, LR76]. Aeration [HS72]. aerial [CH76]. aerial/acoustic [CH76]. Africa [CS74, Llo71, Sze74, Sze76, SST77]. African [Hen74, Ile71a]. After [HH71b, Loc73]. Age [HV73, Loc72, BDC78, Fro78, Hil78, Loc74c, Du 77]. Ageing [MT70]. ages [HH76a]. aid [Lin73, WBH77]. air [KSL74]. Alaskan [PPN78]. alga [TM78]. algae [TM79]. algue [TM78]. algues [TM79]. alimentaire [Cap77, Chr76, TM78, TM79]. Alkalinity [Mor70]. along [Sze74, Sze76]. American [Dow77b, Dow78]. Ammodytes [Pet77, Rea72]. Ammonia [GJ72, GJ74b, Ham77]. ammonium [LT78]. Amounts [CS72]. Amphipodes [RG70]. Amphipods [RG70]. amplitude [Jon73]. anal [Fah79]. Analyse [Lau77]. analyses [SS75]. Analysis [ABL73, Iba72, TL72, Ull77, GJ74b, LT78, Pet79, TLC76]. analytical [RW74]. anchoita [dCdV71]. Anchovy [dCdV71, Hen74]. Anglian [Ile71b]. Angola [dCR71]. Angolan [De 74, dCR73]. Anisakis [Hen74, vB71]. Annual [Loc72, Pla71]. Antillean [Mun74]. Apparatus [Jon72]. appareil [NN73]. Application [CP72, Rea72]. Appraisal [Woo71]. approach [Sze74]. appropriate [Høj78]. Aquaria [JH72, Muž77]. aquatic [vOS76]. aquila [Chr76]. Arctic [GJ74a]. area [BMPC73, Eng76, Hol75b, Sze76, Com74]. Argentine [dCdV71]. Artemia [Jon72, TM79]. artificial [Han74, PPN78]. Ascaroidea [Hen74]. Aspects [Ile71a]. assessment [CH76, Mau77, Smi78]. Assimilation [For72]. Atlantic [BJB71, Dic72, GR74, Gib78, Har76, HH79, KSL74, LMF78, LR76, PR77, WTS73]. atomic [Mor73]. Attachment [NN72]. Attenuation [Wes72]. attractants [Paw77]. Attraction [Han74]. August [Cus74a]. AutoAnalyzer(R) [TLC76]. Automatic [Adr72, GJ72, LT78, vOS76]. automatique [TLC76, LT78]. availability [Dow77a]. avoidance [Han74].

B [Ste75a]. Back [Rea72]. Back-Calculation [Rea72]. Bacteria [Paa71]. Bacteria-Free [Paa71]. Bacterial [Sor72]. bait [Han74, dCR73]. bait-boat [dCR73]. baited [Mil78]. baits [CS78]. Balaenoptera [Loc72, Loc74c]. balanoides [HW75]. Balanus [HW75]. Baltic [Jer75, Len73]. Bank [BMPC73]. barnacle [HW75]. based [Mor73]. Basket [EZKS70]. Bass [Sti72]. batis [Du 77]. Bay [BL71, OKN73, Pla71, Sti72]. be [NN73]. beat [Jon73]. behavior [TM78, TM79]. Behaviour [BLV70, Bur70, Sim71, Muž77, RWHW74, SC74]. Benesch [GJ74b]. benthique [Sar73]. Berend [Kor79]. Bertalanffy [CP72, Loc74b]. Between [Bro71, Dic72, Wie72, Bra79a, GS74, GMD77, Ham77, Hou77, JH78, LT78, Mun74]. Bight [Tal77a]. Binding [For72]. Biochemical [GW71, Com74, HW75]. biological [Pla76, dV76]. Biology [CP72, WW77]. Bjerkan [Rol70]. Black [Sor72]. bladder [KSL74]. Blue [SB71]. boat [Loc74a, dCR73]. Body [Loc76]. Bolinopsis [Len73]. Books [Ano70, Ano71a]. borealis [Loc74c]. Bottom [BBvE71]. Bottom-Trawl [BBvE71]. bovinus [Cap77]. box [Beu74]. Brachionus [How73]. brachyura [Hol72, HT74]. breeding [His75]. Brevoortia [KSL74]. Britain [Bow74, How73]. British [HT74, Hol75a]. Bugt [Pet77]. bull [Cap77].

C [NN72]. côtes [Cap77, Chr76]. caesium [RW74]. caesium-137 [RW74]. calanoid [TM78]. Calanoida [TM79]. calanoides [TM78]. Calanus [Wib76b]. Calcium [Mor70]. Calculation [Rea72, Hol75b]. California [Smi78]. Callionymus [Boe71]. can [NN73]. Canada [Gib78]. Cancer [Ben74, CS78, Mil78]. Cape [BBvE71, CS74]. capensis [BBvE71, Hen74]. captivity [His75, JH78]. Capture [CP72, HH71b]. Carcinus [WD72]. Carruthers [Ram75]. Castelnau [BBvE71]. Catch [Ste75b, De 74, JS74b, Mun74, dCR73, Sis74]. catch-rate [JS74b]. catches [Ben74, CS78, Kim78]. Caudal [Fah72]. caused [Dar74]. cavity [KSL74]. cells [MLG78]. Celtic [Cor74, Du 77]. Celtique [Du 77]. centra [HV73]. central [Las79, WBH77]. Centropages [TM78, TM79]. certain [Gru74b]. Cetacea [BJB71]. chain [JS74b]. Chamber [Thr70b]. Change [Fah72, Fah79]. Changes [GW71, dCR71, Cor74, De 74, Gib78, Jer75, Loc73, Sze76, dV76]. Changing [HSTM71]. channel [DLMS78, HH76b, Hou77]. Character [WBP72, Pos74]. charr [Fro78, KF78]. Chemical [Mor70, Paw77]. chez [BDC78, Du 75]. chimiques [Mar74]. Chlamys [BG75]. Chlorinity [CD70, GW72]. Chlorophyll [Lor71, MS76]. chrome [Mar74]. chromium [Mar74]. Cichlid [Ile71a]. Ciliate [PH71]. circulation [Doo74]. Classification [Jer77]. clausi [TM78]. clavata [Hol72, HV73, HT74, Hol75a]. climatic [Dow77a]. Clupea [BMPC73, Cor74, MT70, WBP72]. clupeoid [NW78]. Clyde [Com74]. Coalfish [Wal71]. Coast [Sze74, HG78, NL78, Sze76, SST77, Llo71, dCR71]. coasts [Cap77, Chr76]. cobalt [Mar74]. Cod [BBvE71, GW70, SG72, Wal71, AH78, Bra79a, GJ74a, JS74a, Jon74, Jon78, LR76, Lov74, Paw77, Pla76, WW77, NB71]. Cod-Ends [BBvE71]. codends [Kim78]. Codworm [Pla76]. Coefficient [Du 75, ABL73]. coefficients [San77]. Cohort [Ull77]. cohortes [BDC78]. cohorts [AC75, BDC78]. colorimetric [GJ74b]. Colour [HJK77, Lov74]. Commercial [SS71, Dow77a, Wib76b]. Committee [Ste74a]. Common [Bro71]. comparaison [LT78]. comparative [LR76, RG70]. compared [Beu74]. Comparées [RG70]. comparison [LT78, MS76]. Component [Iba72]. Comportement [TM78, TM79]. Composantes [Iba72]. Composition [Mor70, Sti72, HW75, Mor73]. Computations [Cus73a, Pet79]. Computer [Sta71, HH74a]. concept [Wib76a]. concerning [Doo74]. Condition [Bro71, GW71, Du 75, GS74, GMD77, WM78, Du 75]. Conditions [GW71]. Conical [RG70]. Coniques [RG70]. conjunction [WBH77]. Considérations [Mar74, Mar74]. Constants [Sta71]. consumption [Jon78]. Content [Bro71, GW71, Com74, GS74, HH76a, WM78]. Continuity [Lor71]. continuous [LM74]. conversion [WS74]. copepod [Dar77, MLG78, Gue70]. Copepoda [GMD77, TM79, Van74]. Copépode [Gue70, Sar73]. Copépodes [RG70, TM78]. Copepods [RG70, TM78]. copper [Hea75]. counter [vOS76]. Counts [Loc72]. Crab [WD72, CS78, Mil79]. crabs [Ben74, Mau77]. Crassostrea [LM74]. crawfish [PR77]. Creel [CS78]. critical [GJ74b, Wib76a]. Critique [MT70]. croissance [Du 77]. Cross [WBP72]. crowding [Fah78]. Crustacé [Gue70]. Crustacea [BDC78, Gue70, OKN73]. Ctenolabrus [Hil78]. Ctenophore [Fra70, Len73]. Cultural [Ste74a]. Culture [Com72, Bow74, How73]. cultured [Bow74]. Cultures [HS72, Paa71, vOS76]. cumulative [Hol75b]. Cunha [PR77]. current [Smi78, SC74]. Curve [CP72, DC79, Eng76, Har75, Hol75b]. cycles [Dow77a, Dow78]. Cyclopoidea [GMD77]. cypris [HW75].

d’âge [BDC78]. Dab [Sim71]. Damage [Dar74]. d’analyse [LT78]. Danish [HH71b]. Danois [Let70]. dans [LT78, Mar74, TLC76]. dark [Pet79]. d’Artemia [TM79]. Data [CP72, Hol72, Iba72, De 74]. day [RWHW74]. Daylight [Høj78]. Decapoda [OKN73]. decay [AC75]. December [Tie76]. definition [BDC78, BDC78]. degree [Pla76]. Delayed [Ric71]. Demand [Adr72]. densities [PPN78]. density [Har75, SM76]. density-dependent [Har75]. Dependence [Cus71a]. dependent [Har75]. Depth [Dic72]. Derived [dV70]. describe [Loc74b]. Description [NN73, NN73]. Design [Adr72]. Detector [SdG71]. Determination [CS72, GJ72, Fro78, GJ74b, MS76, RW74, LT78]. Determined [Hol72, HH79]. determining [Loc74c]. deux [LT78, TM78, TM79]. developing [FO77, Fah78, Fah79, Han74]. Development [GW70, Las72, SB71, BD74, HG78, LR76]. Device [Ric71, BvEH74, NN73, Tan76]. Dicentrarchus [Sti72]. Diclidophora [Smi72]. Diel [Sim71]. diet [Cap77, Chr76]. Dietrich [Lee74]. Diets [For72]. Different [Cus71a, RG70, CS78, HH76a, Lov74]. Différentes [RG70]. Dimensions [RG70]. dioica [Las72]. Direct [GJ72]. discontinuous [LM74]. Disko [Pet77]. dispersal [Tal77a]. Dissolved [BLV70, Bur70]. Distribution [Leb72, Lor71, BMPC73, HD76]. Diurnal [SdG71]. divergent [San77]. Divers [HH71b, SC74]. Does [FO77]. Données [Iba72]. door [JS74b]. door-to-door [JS74b]. Dosage [TLC76, LT78]. Dr. [Ram75]. dredged [DLMS78]. dredging [Dar74]. Drifter [RR72, HH74a, JWA73]. drifters [Dic76]. dry [GS74]. dubius [Pet77]. Due [Wes72]. d’un [NN73, Sar73]. During [BJB71, BLV70, Bur70, Ile71b, SS71, Pet77]. dynamics [Cla74].

eagle [Chr76]. Ear [Loc72, Loc74c]. East [Ile71b]. eastern [HH79, Pos72, Har76]. eau [LT78, Mar74, TLC76]. Echo [Har76, RWHW74]. Echo-traces [Har76]. Ecological [Iba72, Ile71a]. Écologie [Iba72]. Écologiques [Iba72]. Ecology [Fra70, Iba72]. Edouard [Let70]. Educational [Ste74a]. edulis [BGDS78, Dar74, DE76, Dar77, Dav74, GS74, GMD77, HS72, HH76a, How73, MLG78, PPN78, WS74]. Edwards [Hep72]. eel [Pet77]. Effect [Wal71, ABL73, Har75, Hen74, JS74b, JS74a, LM74]. Effects [Dow77a, Dow78, For72, LR76, AH78, Ben74, Cla74, LdBD75, MLG78, Ste74c]. Efficiences [RG70]. Efficiencies [RG70]. efficiency [Beu74, WS74]. efficient [Wei78]. effort [De 74, dCR73, SM76]. Egg [HRH71, Las72, WBH77]. Eggs [SB71, JUH75, Tal77a]. Egypt [Mes77].

Einar [Ruu71]. electric [RM73, Ste74c, SC74]. electrical [Ste75b, Ste77]. electrified [BD74]. elimination [Jon74]. embryo [LR76]. embryonic [HG78]. Ends [BBvE71]. Energy [Bro71, CL78]. England [GMD77, NL78, Sis74]. English [HH76b, Hou77]. Engraulis [Hen74]. enriched [Mar74]. enrichie [Mar74]. entre [LT78]. Entry [Mil78, Mil79]. environmental [Gru74b, SH79]. Enzyme [Dar77]. equation [Cus73a, Loc74b]. equilibrium [Sis74]. Erratum [Ano71b, Ano74a, Ano77a, His76]. Error [ABL73, Wie72]. errors [Ull77]. escapement [Mil79, Mun74]. esmarkii [Smi72]. Establishing [Paa71]. esterase [TM74]. Estimate [Sta71, Gru74a]. Estimates [Jon78]. estimating [Kim78]. Estimation [San77, Eng76, Lau77]. Estuarine [BLV70, Bur70, Dow77a]. Estuary [Bur70, LMF78]. Étude [Cap77, Chr76, Du 75, Sar73]. European [Dow77b, Dow78, Sti72]. Evaluating [De 74]. Evidence [dV70]. examination [DLMS78]. Example [Har70]. Exercise [Wal71]. exhibit [San77]. Experimental [Adr72]. Experiments [DE76, HA71, SdG71, WD72, WS74, dV70]. exploitation [Wib76b]. exploited [Cla74, San77, SH79]. Extinction [Lor72]. extraction [Dar77, Hea75].

F [Len73]. Factor [Bro71]. Factors [RR72, dV70, Gru74b]. Farming [PJ71, PH71]. Fast [Sze76]. feasibility [JUH75]. February [Stu78, Tie76]. Fecundity [BH74, HG78, Las72, HH74b, Hol75a]. Feeders [Adr72]. Feeding [NB71, LM74, TM78, TM79]. female [CL78]. fer [Mar74]. ferrocyanide [Hea75]. ferromagnetic [KSL74]. Fibres [GW70, Wal71, WBP72]. Field [Dic76, Wie72]. fields [RM73, Ste74c, SC74]. Filets [RG70]. Filtering [Saw72]. Fin [Fah72, Loc72, Fah79]. finmarchicus [Wib76b]. Firth [CL78]. Fischer [Sar73]. Fish [Adr72, DVB77, NN72, PJ71, PH71, SdG71, Wes72, Bow74, Cla74, CH76, De 74, Gul77, How73, Loc74b, Mun74, RM78, RWHW74, RM73, San77, Smi78, SC74, WBH77]. Fisheries [CP72, Hem72, Ric71, Wes72, De 74]. Fishery [Gul71b, Wib76b, OKN73, dCR73, Sis74]. Fishes [Cus71a, Ile71a, FO77, Wib76a]. Fishing [SM76, Sta71, SS71, Hou77, Ste75b]. Fitting [Sta71]. fixation [Fah76, Pet79]. flagellates [Paa71]. flatfish [How73, LdBD75, Ste77]. flesh [Com74]. flounder [Sis74]. Fluctuations [Dow72, HH71a]. Flume [HA71]. fluorescence [MS76]. Fol [Las72]. food [HT74, Jon74, Jon78, JH78, Las79, WS74]. foods [How73]. formalin [Loc73, LdBD75]. formes [Mar74]. forms [Mar74]. Fowler [HT74]. fractionating [NN73]. fractionner [NN73]. fragmentary [De 74]. Franca [BBvE71]. Frede [Sme78]. Free [Paa71, Ham77]. French [BDC78, Cap77, Chr76, Du 75, Du 77, Gue70, Iba72, LT78, Mar74, NN73, RG70, TM78, TM79]. frequency [Jon73, SS75]. Friedrich [Tie76]. fulvus [Sar73]. Fundulus [Fah72, Fah76, Fah78, Fah79]. Further [JH78, LdBD75].

gadoids [JH78]. Gadus [GW70, JS74a, Jon74, Jon78, LR76, Lov74, NB71, Paw77, Wal71]. Galeorhinus [HH79]. galeus [HH79]. galilaea [EZKS70]. gammarus [Ben74, HG78, NL78]. Gas [SG72]. general [DC79]. generator [SC74]. genotypes [TM74]. Geoffroy [Cap77]. Georges [BMPC73]. Georgi [Stu78]. gigas [LM74]. glycogen [GS74, HH76a]. Godhavn [Pet77]. grab [Beu74, Pet77]. graphic [AC75]. Greenland [Pla76]. grounds [Lov74]. Group [Ste74a, Loc73, LdBD75]. Groups [Cus71a, Kir74]. Growth [BL71, CP72, For72, GW70, Hep72, Hol72, Ile71a, Ile71b, JH72, Las72, Mau77, Rea72, AC75, BMPC73, DE76, Du 77, Fro78, His75, JH78, LM74, Loc74b, PPN78, PR77, San77, WS74, Bow74]. Gulf [BMPC73, Dow72, Loc74a, MT70]. Günter [Lee74]. Günther [Fro78, KF78].

Haddock [HH71b, JH72, CL78, Jon74, Jon78, LR76]. Hake [BBvE71, CS74]. harengus [BMPC73, Cor74, MT70, WBP72]. Hatchery [GW71, HS72, LMF78, LM74]. hatchery-reared [LMF78]. Hatching [Jon72]. Havinga [Kor79]. Havn [Pet77]. Hela [Voi78]. hemirhamphi [NN72]. Hemirhamphus [NN72]. Henle [HT74]. Henry [LPR70]. Hermann [Sme78]. Herring [Gra78, Ile71b, MT70, WBP72, vB71, BMPC73, BH74, Cor74, Pos74]. Herring-worm [vB71]. Heterocheilidae [vB71]. High [HA71, SC74]. High-speed [HA71]. Hilaire [Cap77]. histochemistry [MLG78]. Histopathological [MLG78]. History [Ile71b, OKN73]. Homarus [Ben74, Hep72, HG78, NL78]. Host [NN72]. Hubbs [dCdV71]. Human [RR72]. Hydrocarbons [CS72]. Hydrodynamics [HA71]. Hydrogen [Sor72]. Hydrostatic [SG72]. Hypotheses [Doo74].

Iceland [GR74, HH74b, Pla76]. Ices [dV70]. Ichthyoneuston [Har70]. Identification [SB71, JUH75, NW78, WBH77]. II [Bur70, RG70, TLC76]. III [Ile71b]. Ilmo [Voi78]. immersion [Ben74]. immunochemical [JUH75]. implications [GJ74a, SH79]. Importance [HS72]. increase [Mes77]. Index [Ano71c, Ano72, Ano74b, Ano76, Ano77b, Ano79, GW71, GS74, GMD77, WM78]. India [Bur70]. Indian [SSS74]. Indicator [CD70, PJ71, HJK77, Pla76]. induced [MLG78]. Inductive [SdG71]. Inexpensive [Jon72, vOS76]. infestation [BGDS78, GMD77]. Infestations [PH71, Van74]. Influence [Fah72, RR72, dCdV71, FO77, Fah78, Fah79, Gru74b, dCR73]. infundibulum [Len73]. ingress [Mun74]. injected [KSL74]. intake [JH78]. intensity [SM76]. International [PSLC76]. Interprétation [Iba72, Iba72]. intestinal [MLG78]. intestinalis [BGDS78, Dar77, GMD77, MLG78]. Invertebrates [Saw72, Han74]. Investigation [Loc74c, Wie72, Ste74c, SST77]. Investigations [Bur70, JH72, BLV70]. ionic [Mor73]. Ireland [Gra78]. iron [Mar74]. irradiance [Jer77]. Isle [BL71]. IX [Bow74].

J [Ram75]. January [Cus74a]. Japan [OKN73]. Jasus [PR77]. Johan [Bra71]. Johannes [Nie76]. John [Ste75a]. Joint [Ste74a]. June [Ste75a]. Juvenile [Har70, Gib78].

kept [GS74]. key [Hil78]. Kirtisinghe [NN72]. Knudsen [GW72]. Krey [Nie76].

L [Com72, Gru74b, HG78, Hil78, HH79, HW75, JH72, Paw77, WS74]. L. [BDC78, BGDS78, Ben74, BMPC73, Bro71, CL78, CS78, Chr76, Com74, Cor74, Dar74, DE76, Dar77, Dav74, GW71, GS74, GW70, Gru74a, HS72, HH71b, HH74b, HM76, HH76a, How73, Jon72, JS74b, JAWS79, Las79, LR76, Loc73, Lov74, MT70, NB71, Pos72, Rea72, Sti72, Wal71, WBP72, WJA78, WD72, dV70, dV76, vB71]. Labidocera [Gue70]. Laboratory [Ste74a, Bow74, GS74]. labrax [Sti72]. Lafont [HT74]. l’Aigle [Chr76]. Lake [Mes77]. l’Analyse [Iba72]. landings [Dow77b, Dow78]. Large [Jon72, Loc76]. Large-Scale [Jon72]. larva [HW75]. Larvae [HS72, NB71, SB71, How73, NW78, Tal77a]. Larval [Com72, Wib76a, BMPC73, Hen74, How73, Las79, NL78, WW77]. Latitudinal [Leb72]. Lawrence [MT70]. Laxey [BL71]. Layer [Loc72]. Layers [Dic72]. Laying [HRH71]. Lea [Ruu71]. Length [Wie72, Gru74b, Hil78, KF78, Loc73, LdBD75]. Lernaeenicus [NN72]. Lernaeocera [Van74]. Life [Ile71b]. Light [Lor72]. lignin [PH74]. limitations [Ull77]. Linnaeus [HT74, NL78]. lobster [Dow77b, HG78, NL78]. Lobsters [Hep72, Ben74, Dow78, Mau77]. Locomotion [SdG71]. Logistic [Kim78]. Long [Jer75, dCR71]. Long-Term [dCR71]. Loss [KSL74]. low [JS74a]. Lowestoft [HA71, HSTM71]. lucens [OKN73].

m [Pet77, Thr70a, Thr70b]. Mêmes [RG70]. Mackerel [Pos72, LH74]. Macroplankton [Lin73, HD76]. maenas [WD72]. Maille [RG70]. Mailles [RG70]. Maine [Dow72, BMPC73, LMF78]. Maintained [GW71]. majalis [Fah72, Fah76, Fah78, Fah79]. male [CL78]. Man [BL71]. Management [Gul71b, GJ74a, SH79]. manganese [Mar74, Mar74]. Mangelsdorf [GJ74b]. mantis [BDC78]. mapping [Smi78]. March [Ram75, Ste75a]. Margaret [Pla71]. marina [vB71]. Marine [Bow74, How73, PJ71, PH71, Saw72, Dow77a, Han74, PH74, Ste74a, SC74, Wib76a]. marketable [Bow74]. Martni [dCdV71]. Matamek [Gib78]. material [SD76]. matière [Mar74]. matter [Mar74]. Maturity [Loc72]. maximum [Gru74b]. maximus [Com72, Com74, Gru74a, Gru74b, Las79]. May [Stu78]. means [Loc74c]. measurements [Høj78]. Mechanism [HSTM71]. mechanized [Dar74]. media [TM78, TM79]. Melanogrammus [CL78, HH71b, JH72, Jon74, Jon78, LR76]. Menai [DE76, Hep72, JS70]. Mengraulis [dCdV71]. menhaden [KSL74]. mer [Chr76, LT78, Mar74, NN73, TLC76, Du 77, Iba72]. Mercenaria [WD72].

Mercury [GR74, WTS73]. Merlangius [HH74b, His75, HM76, JH72, Jon74]. merlangus [HH74b, His75, HM76, JH72, Jon74]. merluccius [BBvE71, BBvE71]. Mesh [RG70]. Meshes [RG70]. Method [GJ72, Paa71, Sta71, Dav74, GJ74b, Hol75b, Kir74, RW74, Wei78]. méthodes [LT78]. Methods [For72, TL72, GJ74b, JUH75, LT78, Ste74a, TLC76]. Meyer [Tie76]. Meyer-Waarden [Tie76]. Micromesistius [SB71]. Microscopes [Thr70b]. migration [HH76b, JAWS79, LMF78, RM73, Wei78]. Migrations [Gue70, LH74, HH79]. mile [RWHW74]. milieux [TM78, TM79]. Milne [Hep72]. Milne-Edwards [Hep72]. Minch [HH74b]. Mixing [BLV70, Bur70, Pla76]. mode [DLMS78, Mun74]. Model [Leb72, Kim78]. Modern [GW72]. modified [Loc74a]. Mohr [GW72]. Molluscan [Dow72]. monitor [Sze74]. monitored [Sze73]. monitoring [Dav74]. Monogenea [Smi72]. montagui [Hol72, HT74]. Moray [CL78]. morhua [GW70, JS74a, Jon74, Jon78, LR76, Lov74, NB71, Paw77, Wal71]. morphological [Fah76]. mortality [ABL73, Cus75a, Gru74a, Har75, LR76, SM76, San77, Wib76a]. Moulting [Hep72, PR77]. Mourine [Cap77]. movements [JWA73, LMF78, WJA78]. moving [Paw77]. Muller [How73, HT74, Len73]. multichannel [Sze74]. Multipurpose [HSTM71]. Muscle [GW70, TM74, Wal71, WBP72]. musculature [WW77]. mussel [How73]. mussels [Dar74, DE76, Dar77, Dav74]. Myliobatidae [Cap77]. Myliobatis [Chr76]. Mytilicola [BGDS78, Dar77, GMD77, MLG78]. Mytilus [BGDS78, Dar74, DE76, Dar77, Dav74, GMD77, How73, MLG78, PPN78, WD72].

N [Ram75, RG70]. naevus [Du 75, Du 77, HT74]. Naples [Sti72]. Nations [Ste74a]. natural [ABL73, Cus75a, Gru74a, Høj78]. Nature [NN72]. nauplii [TM79]. N^{o} [RG70]. neevus [Du 75, Du 77]. Nematoda [Hen74, vB71]. nematodes [WW77]. Nephrops [Ste74c, SS71]. Net [HSTM71, HH71b, Wie72]. Net-Changing [HSTM71]. Netherlands [BD74]. Nets [RG70]. neutral [Loc73, LdBD75]. Nikolsky [Stu78]. nilotica [EZKS70]. Nils [Sah71]. Nilsson [Smi72]. Nitrogen [Bro71]. noise [SH79]. Non [CS72, Thr70a, Mar74]. Non-Polar [CS72]. Non-Toxic [Thr70a]. Nord [Iba72]. North [Boe71, DE76, Hep72, Llo71, Pos72, Smi72, Tal77a, BH74, Dic72, Doo74, Har76, HH71a, HG78, HH74b, HM76, HH79, Iba72, Ile71b, JAWS79, Las72, Las79, LH74, Pos72, Smi72, WBP72, WBH77, WTS73, dV76]. North-Eastern [Pos72, Har76]. North-West [Llo71]. northeast [NL78, GJ74a]. northern [Doo74, HM76, LH74, Pos72, Smi72]. northwest [SST77]. norvegicus [Ste74c]. Norway [Dow78, Smi72, Wib76b]. Note [Gue70, HH71a, Eng76, GS74, HD76, Gue70]. Notes [GJ74b, RM78, TL72, TLC76, vB71]. Nova [Pla71]. November [Lee74, Ram75]. nucleus [Pos74]. Number [Fah72, Leb72, dCdV71, Bra79a, Fah76, FO77, Fah78, Fah79, Kir74].

Numbers [WBP72]. nutrient [TM78, TM79]. nutritifs [TM78, TM79]. NW [Sze74, Sze76]. Nycthemeral [Gue70]. Nycthémérales [Gue70].

O. [Len73]. Obituaries [Ste75a]. Obituary [Stu78]. Observations [CS74, HH71b, Bla75, JH78, LdBD75]. Occurrence [Boe71, Smi72, Len73, NL78, Pet77]. Ocean [Dic72, Lor72, Ham77, HJK77, Sze73, SSS74]. Oceanic [Ste74a, RM73]. Oceanographic [PSLC76, Gra78]. ocellata [TM74]. October [Lee74]. Off [Llo71, CS74, Gra78, HG78, NL78, RM78, dCR71, SST77]. ogives [Kim78]. Oikopleura [Las72]. Oil [CS72]. open [LMF78, Loc74a, WJA78]. open-water [LMF78]. operation [Mun74]. opercularis [BG75]. optical [Jer75]. ordinate [Hol75b]. Orfordness [RM78]. organic [Mar74, WM78]. organique [Mar74]. Organisation [Ste74a]. organisms [vOS76]. Orientating [dV70]. Orientation [dV70]. orthophosphates [TLC76]. Ostrea [GS74, HS72, HH76a, WS74]. Ostreaedulis [GW71]. Otolith [Rea72, Bla75, Pos74]. Otoliths [MT70]. otter [JS74b]. overlap [Kim78]. Oxidation [Sor72]. oxygen [PSLC76]. oyster [WM78]. Oysters [GW71, GS74, HH76a].

pêche [Lau77]. pagurus [Ben74, CS78]. paradoxus [BBvE71]. parameters [dV76]. parasitic [BGDS78, Dar77]. parasitica [Van74]. Parent [Cus71a]. Pattern [Rea72, SD76]. Paul [Rol70, Tie76]. Pecten [Com72, Com74, Gru74a, Gru74b]. pelagic [CH76, Smi78]. Penaeidea [OKN73]. Penobscot [LMF78]. performance [AH78]. Period [HH71a, Pos72, Jer75, Wib76a]. Petersen [Cus76a]. Phenosafranin [CD70]. Photography [Hem72]. photosynthetic [HJK77, Høj78]. Phototactic [Paa71]. physalus [Loc72]. physical [AH78]. physico [Mar74]. physico-chimiques [Mar74]. physicochemical [Mar74]. Physiological [BGDS78]. physostome [SS75]. Phytoplankton [JS70, Leb72, Lor72, PPN78, Pla71, Woo71, MS76]. Pike [Har70]. Pilchard [Sim71, TM74]. pilchardus [Muž77]. pileus [Fra70]. Pisces [Cap77]. Plaice [BL71, NB71, Sim71, dV70, AH78, Bla75, Bow74, Cus75a, HH76b, JS74b, JAWS79, Loc73, Tal77a, WJA78]. Plancton [RG70, NN73]. Planctonique [Iba72]. Plankton [HA71, Iba72, RG70, Thr70b, BvEH74, Kir74, Loc74a, NN73, WBH77, WTS73]. platessa [JS74b, JAWS79, Loc73, NB71, WJA78, dV70]. Pleurobrachia [Fra70]. Pleuronectes [JS74b, JAWS79, Loc73, NB71, WJA78, dV70]. pleuronectid [JUH75]. Pleuronectidae [Ste77]. plicatilis [How73]. Plug [Loc72, Loc74c]. poids [Du 75]. Polar [CS72]. polymorphic [BG75]. Population [ABL73, HM76, Sor72, Ull77, AC75, Gib78, Gru74a, dCR73]. Populations [Sim71, Cla74, San77, SH79]. Possible [Cla74, Han74, Pla76]. pot [Ben74]. Pour [RG70]. Pout [Smi72]. poutassou [SB71]. Power [Sta71, Hou77, SM76]. praise [Cus76a]. Prawn [For72]. precise [RW74]. Precision [Smi78]. Predation [WD72]. Prediction [RM73]. Preliminary [Bur70]. Preparation [Hea75]. presence [TM79, TM79]. present [Cus78, OKN73]. preservation [Loc73, LdBD75]. Pressure [SG72]. Primary [Llo71, Pet79]. Principal [Iba72]. Principales [Iba72]. Problem [PH71, Mes77]. procedure [PH74]. Processes [Sor72]. Production [Llo71, Pla71, Pet79]. productus [Mil78]. Programme [Sta71]. properties [Jer75]. proposed [CH76, Mes77]. Proximate [Sti72]. Pteromylaeus [Cap77]. puissances [Lau77]. pulse [SC74].

Qarun [Mes77]. quanta [Jer77]. Quebec [Gib78].

R. [Cus74a, Hol72]. Races [BJB71]. racial [Pos74]. radio [Sar73]. Raja [Du 75, Du 77, Hol72, HV73, HT74, Hol75a]. Rajidae [Du 75]. rapid [Kir74, PH74]. Rate [HS72, HRH71, HG78, JS74b, Jon74]. Rates [Hol72, San77]. Ray [HRH71, Cap77]. Rays [Fah72, Fah79]. reared [LMF78, LM74]. Rearing [vB71]. Recapture [CP72]. Received [Ano70, Ano71a]. recirculation [WS74]. recognition [SD76]. record [RWHW74]. Recording [Saw72]. Recoveries [RR72]. Recruitment [Cus71a, DC79, GJ74a, Har75, Sis74]. reduced [Mil79]. reducing [dCR73]. régime [Cap77, Chr76]. Reineck [Beu74]. relaid [DE76]. Relation [SG72, Du 75, Gra78, JH78, RM78, RM73]. Relationship [Dic72, Dow77b, Ham77, Wie72, Bra79a, GS74, GJ74a, GMD77, KF78]. Relationships [Bro71, Mun74]. relative [Dow77a]. Release [Ric71]. remarkable [Van74]. Reply [Nai74]. report [Ste74a]. Research [Hem72, Ste74a]. reserves [CL78]. resonance [SS75]. Respectively [Wal71]. response [Ste77]. responses [BGDS78, Paw77]. results [Ste74a]. reticulatus [Boe71]. Returned [SS71]. returns [HH74a]. review [AC75, GJ74b]. Reviews [Arn73, Bak78, Bea71, Bir74, Bow76, Bra72a, Bra72b, Bra79b, Bur74, Cri71, Cul74, Cur71, Cur72, Cus71b, Cus71c, Cus72a, Cus72b, Cus73b, Cus74b, Cus74c, Cus74d, Cus75b, Cus76b, Ell72, Fle79, Fra71, Gar76, Gar77, Gou74, Gul71a, Gul72, Gul79, Haw78, Hem71, Her74, Hol71, Hor75, Jam71, Jer71, Jon77, Kas71, Kin73, Lar71, Let71, Lis74, Lis76, Lon74a, Lon72, Lon74b, Lon74c, Lyt78, McC74, Nic72, Pau71, Paw76, Pér73, Por77, Por78, Pre73, Pre74, Pre77a, Pre77b, Pre78, Pur74, Ruf74, Sch72, Sco73, Sig71, Sme71, Smi76, Sor71, Ste74b, Tai74, Tal77b, Wal78a, Wal78b, Wer77]. rhodamine [TB75]. Risso [SB71]. River [Gib78, LMF78]. role [Han74]. Rollefsen [Sæt77]. Rosén [Sah71]. rotifer [How73]. rupestris [Hil78]. Ruud [Bra71].

S [Cus74a]. safe [SC74, Tan76]. Saint [Cap77]. Saint-Hilaire [Cap77]. salar [LMF78]. salina [Jon72]. Salinity [PJ71, FO77, LR76, Mes77]. Salmo [LMF78]. salmon [Gib78, LMF78]. Salmonid [AR77]. Salvelinus [Fro78, KF78]. Same [RG70]. Sampler [HSTM71, Thr70a, Beu74, Loc74a]. Samplers [HA71]. samples [BvEH74, Pet77]. Sampling [Wie72]. sand [Pet77]. sarcoplasm [BG75]. Sardina [Muž77]. sardines [Muž77].

Sardinops [TM74]. Satellites [Sze72, Sze73]. Saturation [Mil79]. saurus [Har70]. Saury [Har70]. Scale [Jon72]. scales [Fro78]. Scallop [Com72, Com74, Gru74a, Gru74b]. scan [DLMS78]. scanning [Dic76, DLMS78, JWA73, Jon73]. Scattering [Dic72]. school [DVB77]. schooling [Cla74]. Science [Gul71b]. Scientific [Ste74a]. Scomber [Pos72]. Scomberesox [Har70]. scombrus [Pos72]. Scophthalmus [Las79]. Scotia [Pla71]. Scotland [Smi72]. Scottish [Fra70, WW77]. screening [PH74]. Sea [CS72, Com74, GW72, GJ72, Leb72, SS71, Chr76, Dow77b, Dow78, GJ74b, Hea75, Høj78, Jer77, Mor73, NN73, PSLC76, RW74, Sze74, SSS74, TLC76, WJA78, Boe71, BH74, Cor74, Doo74, Dow72, Du 77, HH71a, HH74b, HM76, Iba72, Ile71b, JAWS79, Las72, Las79, LH74, Pos72, Smi72, Sor72, Tal77a, WBP72, WBH77, dV76]. sea-surface [Dow78, SSS74]. Seabed [RR72, Dic76, JWA73]. Seasonal [BMPC73, Com74, Rea72, Loc74b]. Seaward [LMF78]. Seawater [Mor70, TL72, LT78, Mar74]. Section [WBP72]. sector [Dic76, DLMS78, JWA73, Jon73]. sector-scanning [Dic76]. Sedimentation [Thr70b, Tan76]. sediments [PH74, TB75]. seed [DE76]. sei [Loc74c]. Seine [HH71b]. Selection [BBvE71, Kim78]. Selective [JAWS79]. Sélectivité [RG70]. Selectivity [EZKS70, RG70, Ste75b]. semiquantitative [PH74]. sensibilité [Sar73]. Sensitive [GJ72]. sensitivity [SH79]. Separation [WBP72]. Sergestes [OKN73]. seule [TM78]. sexes [San77]. Sexual [Loc72]. Shallow [Thr70a]. shape [Har75]. shark [Jon73]. shoaling [CS74, Muž77]. Shore [WD72]. Short [HH71a]. shrimp [BD74]. shrimp-trawl [BD74]. shrimps [Mau77]. side [DLMS78]. side-scan [DLMS78]. Sightings [BJB71]. Significance [Wes72]. Silicon [BLV70, Bur70]. Simple [Paa71, RW74]. Simulation [BDC78, HH74a, BDC78]. single [TM78]. Size [Wie72, Bow74, Du 75, Kir74]. size/weight [Du 75]. Skagerak [LH74]. Skeletal [GW70, Wal71, WBP72]. Skipjack [dCR71]. Skylab [SST77]. Small [CS72, SS71, Thr70b, Loc74a, WM78]. smolts [LMF78]. soak [Mun74]. Sole [Bro71, dV76]. Solea [Bro71, dV76]. solubility [PSLC76]. solution [Mes77]. Some [For72, vB71, Loc76, dV76]. sonar [Cus73a, Dic76, DLMS78, JWA73, Jon73, Smi78]. Sonic [Dic72]. Sorting [NW78, Dar74, Kir74, Lin73]. Sources [Ull77]. South [PR77, CS74, Hen74]. Southampton [BLV70]. Southern [Bur70, Loc72, JAWS79, Loc74c, SSS74, Boe71, Sis74, Tal77a]. southwest [GMD77]. sp [TM79]. sp. [TM79]. space [Sze74]. spat [LM74, WS74, WM78]. spatfall [Dav74]. Spawning [Sim71, HH76b]. Species [HRH71, Leb72, Dow77a, Eng76, Hol75b, Loc76]. species-area [Eng76]. Specific [Dow72]. spectrophotometric [MS76]. speed [HA71, Jon73]. spp [Van74]. Sprat [Sim71]. Squilla [BDC78]. St [MT70]. St. [Pla71]. stability [Gul77, Lov74]. Stage [Ile71b]. stages [NL78]. Starvation [Wal71]. state [Cus78]. status [OKN73]. Stereoscan [Bla75]. Steuer [Dar77, GMD77]. stimulation [Ste77]. Stock [Cus71a, GJ74a, WBP72, Cor74, CH76, DC79, De 74, Har75]. stock-recruitment [DC79]. stocks [Gul77, Pla76]. stomachs [Jon74]. Stomatopoda [BDC78]. Straits [DE76, Hep72, JS70]. strategy [CH76]. stream [JAWS79, Wei78]. streams [RM78]. strength [DVB77]. structure [BDC78]. Studies [Ile71b, Rea72, SS75, Sze72, WW77, BH74, HM76, HJK77, Høj78, HV73, SC74]. Study [JUH75, Woo71, Ste74a, Ste77, Cap77, Chr76, Du 75]. subsampling [BvEH74, HD76]. Successful [vB71]. Sulphide [Sor72]. summary [Ste74a]. summer [HH76a]. supporting [dCR73]. Surface [Lor71, CS74, Dow77b, Dow78, HH74a, Sze73, Sze74, SSS74]. surface-shoaling [CS74]. Suruga [OKN73]. survey [Cus78]. surveys [WBH77]. Survival [HH71b, SS71, DE76]. suspended [SD76, TB75]. Swedish [Hil78]. Swimbladder [SG72, SS75]. swimming [AH78, Jon73]. system [BG75, PPN78, WS74]. systems [Ste75b].

T [Bra71]. T. [EZKS70]. table [Mor73]. Tables [PSLC76]. tag [AH78]. Tagged [HH71b]. Tagging [Hol72, HH79]. tags [KSL74]. Tail [Jon73]. taille [Du 75]. taille/poids [Du 75]. Tait [Ste75a]. taken [JS74b, Pet77]. target [DVB77]. Technique [Saw72, Tan76]. Techniques [TL72, Dic76, GJ74b, NW78, TLC76]. Temperature [Dow72, Fah72, HH71a, dCdV71, Bra79a, Dow77b, Dow78, Fah79, JS74a, LR76, Sze74, Sze76]. temperatures [Sze73, SSS74]. Term [dCR71]. terminology [AR77]. terms [Jer77]. terrestrial [Ham77]. tests [Dic76]. tetracycline [HV73]. Their [For72, Fro78, Han74]. theoretical [HD76]. Three [HRH71, RWHW74]. Thunberg [LM74]. tickler [JS74b]. Tidal [RR72, Wei78, JAWS79, RM78]. Tigriopus [Sar73]. Tilapia [EZKS70]. time [Ben74, Fah76]. Titration [GW72]. Titrations [CD70]. tobianus [Rea72]. tope [HH79]. Total [WBP72, Fah76, Fah78, Ham77]. Tow [Wie72]. Toxic [Thr70a]. traces [Har76]. tracked [JWA73, Jon73, WJA78]. tracking [RM78]. Trans [BJB71]. Trans-Atlantic [BJB71]. Transparency [Dic72]. Transplantation [dV70]. transport [JAWS79, Wei78]. Trap [Ric71]. Traps [EZKS70, Mil78, Mil79, Mun74]. Trawl [BBvE71, HH71b, BD74, JS74b]. trawlers [Hou77]. trawling [WBH77]. Trisopterus [Smi72]. Tristan [PR77]. tristani [PR77]. tuna [dCR73]. Tunisian [Cap77, Chr76]. tunisiennes [Cap77, Chr76]. turbot [Las79]. twenty [RWHW74]. twenty-mile [RWHW74]. twenty-three [RWHW74]. Two [BJB71, Van74, LT78, TM78, TM79]. typicus [TM78, TM79]. tyrannus [KSL74].

uncertainties [ABL73]. Underwater [Hem72]. unfished [Gru74a]. United [Ste74a]. Untagged [HH71b]. upon [Fah78]. Upwelling [Sze72, PPN78, Sze76, SST77]. Use [PJ71, Fro78, Loc74b, Loc74a, MT70, Ric71, SdG71, SC74, Thr70b, WBH77]. used [NN73]. Using [Loc72, CS78, Dic76, DLMS78, HV73, Tan76]. Utermöhl [Tan76]. utilisable [NN73]. utilisation [TLC76].

V [Stu78, Loc74a, NN72]. vachette [Cap77]. valid [Loc74c]. validation [HV73]. Validity [Sze73]. values [Hol75b]. Variability [Sis74, SSS74]. Variation [Mor70, Gra78]. variations [Com74]. Veen [Beu74, Pet77]. Vellar [Bur70]. Version [GW72]. versus [LM74]. Vertebrae [Fah72, dCdV71, Fah76, FO77, Fah78]. vertebral [Bra79a]. VIII [How73]. virens [Wal71]. Virtual [ABL73, Ull77]. vivo [MS76]. Volume [Ano71c, Ano72, Ano74b, Ano76, Ano77b, Ano79]. vulgaris [Hep72].

Waarden [Tie76]. Walb [Har70]. Walbaum [Fah72, Fah76, Fah78, Fah79]. Wales [DE76, Hep72, HG78]. warning [Hea75]. Water [CS72, GW72, GJ72, Thr70a, Bra79a, GJ74b, Hea75, Jer77, LMF78, Mor73, Paw77, PSLC76, RW74, TLC76, BLV70]. Waters [Fra70, Thr70a, Ham77, Hil78, Høj78, HT74, Hol75a, WW77]. Weight [Bro71, Loc73, Du 75, GS74, KF78, LdBD75]. weights [Com74, Loc76, Mor73]. West [CS74, Hen74, Llo71, Gra78, Las79, WBH77, Smi72]. western [Hou77, Len73]. Wet [Bro71]. Whale [Loc72, Loc74c]. whales [Loc76]. which [San77]. White [WBP72]. Whiting [JH72, SB71, HH74b, His75, HM76, Jon74, WW77]. whose [Kim78]. Wild [Thr70b]. willughbii [Fro78, KF78]. Wimpenny [Cus74a]. Wind [RR72]. Windermere [Fro78, KF78]. winter [HH76a, Pet77]. Wire [EZKS70]. Wood [LPR70]. Woodhead [JWA73, RR72]. Working [Ste74a]. worm [vB71].

xanthopterus [NN72].

Yacht [BJB71]. Years [Ile71b, dV70]. Yellowfin [dCR71, dCR73]. yellowtail [Sis74]. yield [DE76].

Zeiss [Thr70b]. zillii [EZKS70]. zinc [Mar74]. zooplankton [Ste74a].

References

Agger:1973:EVP

[ABL73] P. Agger, I. Boëtius, and H. Lassen. Error in the virtual population analysis: the effect of uncertainties in the natural mortality coefficient. ICES Journal of Marine Science, 35(1):93, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

          tronic). URL http://academic.oup.com/icesjms/article/35/
          1/93/769838.

Alverson:1975:GRG

[AC75] Dayton L. Alverson and Michael J. Carney. A graphic review of the growth and decay of population cohorts. ICES Journal of Marine Science, 36(2):133–143, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/36/2/133/746294.

Adron:1972:DAD

[Adr72] J. W. Adron. A design for automatic and demand feeders for experimental fish. ICES Journal of Marine Science, 34(2):300–305, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
34/2/300/595538.

Arnold:1978:PEA

[AH78] G. P. Arnold and B. H. Holford. The physical effects of an acoustic tag on the swimming performance of plaice and cod. ICES Journal of Marine Science, 38(2):189–200, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
2/189/743163.

Anonymous:1970:BR

[Ano70] Anonymous. Books received. ICES Journal of Marine Science, 33(2):300, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/300/800745.

Anonymous:1971:BR

[Ano71a] Anonymous. Books received. ICES Journal of Marine Science, 34(1):157, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/157/602202.

Anonymous:1971:E

[Ano71b] Anonymous. Erratum. ICES Journal of Marine Science, 33(3):514, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
33/3/514/711583.

Anonymous:1971:IV

[Ano71c] Anonymous. Index to volume 33. ICES Journal of Marine Science, 33(3):522–524, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/522/711626.

Anonymous:1972:IV

[Ano72] Anonymous. Index to volume 34. ICES Journal of Marine Science, 34(3):550–552, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/550/758650.

Anonymous:1974:E

[Ano74a] Anonymous. Erratum. ICES Journal of Marine Science, 35(3):379, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/3/379/672892.

Anonymous:1974:IV

[Ano74b] Anonymous. Index to volume 35. ICES Journal of Marine Science, 35(3):380–382, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/380/672896.

Anonymous:1976:IV

[Ano76] Anonymous. Index to volume 36. ICES Journal of Marine Science, 36(3):302, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/302/639301.

Anonymous:1977:E

[Ano77a] Anonymous. Erratum. ICES Journal of Marine Science, 37 (3):323, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/323/637322.

Anonymous:1977:IV

[Ano77b] Anonymous. Index to volume 37. ICES Journal of Marine Science, 37(3):324–326, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/324/637326.

Anonymous:1979:IV

[Ano79] Anonymous. Index to volume 38. ICES Journal of Marine Science, 38(3):351–353, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/351/646641.

Allan:1977:ST

[AR77] I. R. H. Allan and J. A. Ritter. Salmonid terminology. ICES Journal of Marine Science, 37(3):293–299, September 1977. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/37/3/293/
637256.

Arnold:1973:R

[Arn73] G. P. Arnold. Reviews. ICES Journal of Marine Science, 35 (1):104–105, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/104/769594.

Baker:1978:R

[Bak78] C. W. Baker. Reviews. ICES Journal of Marine Science, 38 (2):273, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/273/743272.

Bohl:1971:SCH

[BBvE71] H. Bohl, L. Botha, and T. H. van Eck. Selection of cape hake (Merluccius merluccius capensis Castelnau and Merluccius merluccius paradoxus Franca) by bottom-trawl cod-ends. ICES Journal of Marine Science, 33(3):438–471, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/3/438/711526.

Boonstra:1974:DES

[BD74] G. P. Boonstra and S. J. De Groot. The development of an electrified shrimp-trawl in the Netherlands. ICES Journal of Marine Science, 35(2):165–170, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/2/165/695990.

Badia:1978:SSA

[BDC78] Jacques Badia and Thang Do-Chi. Simulation de la structure d’âge et définition des cohortes chez Squilla mantis (L.) (Crustacea, Stomatopoda). (French) [Simulation of age structure and definition of cohorts in Squilla mantis (L.) (Crustacea, Stomatopoda)]. ICES Journal of Marine Science, 38(1):105–115, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
1/105/630175.

Beamish:1971:R

[Bea71] F. William H. Beamish. Reviews. ICES Journal of Marine Science, 34(1):144–145, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/144/602178.

Bennett:1974:EPI

[Ben74] D. B. Bennett. The effects of pot immersion time on catches of crabs, Cancer pagurus L. and lobsters, Homarus gammarus (L.). ICES Journal of Marine Science, 35(3):332–336, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/332/672818.

Beukema:1974:EVV

[Beu74] J. J. Beukema. The efficiency of the Van Veen grab compared with the Reineck box sampler. ICES Journal of Marine Science, 35(3):319–327, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/319/672810.

Beaumont:1975:PSS

[BG75] A. R. Beaumont and Ll. D. Gruffydd. A polymorphic system in the sarcoplasm of Chlamys opercularis. ICES Journal of Marine Science, 36(2):190–192, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/36/2/190/746388.

Bayne:1978:PRM

[BGDS78] B. L. Bayne, J. M. Gee, J. T. Davey, and C. Scullard. Physiological responses of Mytilus edulis L. to parasitic infestation by Mytilicola intestinalis. ICES Journal of Marine Science, 38(1):12–17, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/12/630182.

Burd:1974:FSN

[BH74] A. C. Burd and G. J. Howlett. Fecundity studies on North Sea herring. ICES Journal of Marine Science, 35(2):107–120, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/2/107/695968.

Birkett:1974:R

[Bir74] Leon Birkett. Reviews. ICES Journal of Marine Science, 36 (1):97, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/97/642688.

Brown:1971:SCD

[BJB71] S. G. Brown, S. J. Jones, and A. Boyde. Sightings of cetacea during two trans-Atlantic yacht races. ICES Journal of Marine Science, 34(1):106–120, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/106/602154.

Bowers:1971:GPL

[BL71] A. B. Bowers and Jong Wha Lee. The growth of plaice in Laxey Bay (Isle of Man). ICES Journal of Marine Science, 34(1):43–50, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
34/1/43/602216.

Blacker:1975:SOP

[Bla75] R. W. Blacker. Stereoscan observations of a plaice otolith. ICES Journal of Marine Science, 36(2):184–187, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/184/746376.

Burton:1970:BDSa

[BLV70] J. D. Burton, P. S. Liss, and V. K. Venugopalan. The behaviour of dissolved silicon during estuarine mixing I. Investigations in Southampton Water. ICES Journal of Marine Science, 33(2):134–140, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/134/800642.

Boyar:1973:SDG

[BMPC73] H. C. Boyar, R. R. Marak, F. E. Perkins, and R. A. Clifford. Seasonal distribution and growth of larval herring (Clupea harengus L.) in the Georges Bank–Gulf of Maine area from 1962 to 1970. ICES Journal of Marine Science, 35(1):36–51, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/35/1/36/769747.

Boer:1971:OCR

[Boe71] Peter Boer. The occurrence of Callionymus reticulatus in the Southern North Sea. ICES Journal of Marine Science, 33(3):506– 509, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/33/3/506/711560.

Bowers:1974:MFC

[Bow74] A. B. Bowers. Marine fish culture in Britain IX. Growth of cultured plaice to marketable size in the laboratory. ICES Journal of Marine Science, 35(2):149–157, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/2/149/695981.

Bowden:1976:R

[Bow76] K. F. Bowden. Reviews. ICES Journal of Marine Science, 36 (3):299–300, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/299/639294.

Braarud:1971:JR

[Bra71] Trygve Braarud. Johan T. Ruud 1903–1970. ICES Journal of Marine Science, 33(3):308–311, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/3/308/711464.

Brander:1972:Ra

[Bra72a] K. M. Brander. Reviews. ICES Journal of Marine Science, 34 (2):317–319, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/317/595550.

Brander:1972:Rb

[Bra72b] K. M. Brander. Reviews. ICES Journal of Marine Science, 34 (3):545–546, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/545/758633.

Brander:1979:RBV

[Bra79a] K. Brander. The relationship between vertebral number and water temperature in cod. ICES Journal of Marine Science, 38 (3):286–292, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/286/646585.

Brander:1979:R

[Bra79b] K. M. Brander. Reviews. ICES Journal of Marine Science, 38 (3):348–350, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/348/646636.

Bromley:1971:RBW

[Bro71] P. J. Bromley. Relationships between the wet weight condition factor and nitrogen energy content in the common sole (Solea solea L.). ICES Journal of Marine Science, 34(1):131–134, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
1/131/602163.

Burton:1970:BDSb

[Bur70] J. D. Burton. The behaviour of dissolved silicon during estuarine mixing II. Preliminary investigations in the Vellar Estuary, Southern India. ICES Journal of Marine Science, 33(2):141–148, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/141/800650.

Burd:1974:R

[Bur74] A. C. Burd. Reviews. ICES Journal of Marine Science, 35 (2):217, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/217/696073.

Burrell:1974:NDS

[BvEH74] V. G. Burrell, Jr., W. A. van Engel, and S. G. Hummel. A new device for subsampling plankton samples. ICES Journal of Marine Science, 35(3):364–366, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/35/3/364/672843.

Capape:1977:EDR

[Cap77] Christian Capapé. Étude du régime alimentaire de la mourine vachette, Pteromylaeus bovinus (Geoffroy Saint-Hilaire, 1817) (Pisces, Myliobatidae) des côtes tunisiennes. (French) [Study of the diet of the bull ray, Pteromylaeus bovinus (Geoffroy Saint-Hilaire, 1817) (Pisces, Myliobatidae) from the Tunisian coasts]. ICES Journal of Marine Science, 37(3):214–220, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
3/214/637204.

Cunningham:1970:PIC

[CD70] C. C. Cunningham and I. W. Duedall. Phenosafranin indicator for chlorinity titrations. ICES Journal of Marine Science, 33 (2):292–293, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/292/800718.

Cram:1976:PAA

[CH76] D. L. Cram and I. Hampton. A proposed aerial/acoustic strategy for pelagic fish stock assessment. ICES Journal of Marine Science, 37(1):91–97, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/1/91/622196.

Christian:1976:EDR

[Chr76] Capapé Christian. Étude du régime alimentaire de l’Aigle de mer, Myliobatis aquila (L., 1758) des côtes tunisiennes. (French) [Study of the diet of the sea eagle, Myliobatis aquila (L., 1758) from the Tunisian coasts]. ICES Journal of Marine Science, 37(1):29–35, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
37/1/29/622178.

Campbell:1978:ERM

[CL78] Susan Campbell and R. M. Love. Energy reserves of male and female haddock (Melanogrammus aeglefinus L.) from the Moray Firth. ICES Journal of Marine Science, 38(1):120–121, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
1/120/630186.

Clark:1974:PES

[Cla74] Colin W. Clark. Possible effects of schooling on the dynamics of exploited fish populations. ICES Journal of Marine Science, 36(1):7–14, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/7/642641.

Comely:1972:LCS

[Com72] C. A. Comely. Larval culture of the scallop Pecten maximus (L). ICES Journal of Marine Science, 34(3):365–378, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/365/758471.

Comely:1974:SVF

[Com74] C. A. Comely. Seasonal variations in the flesh weights and biochemical content of the scallop Pecten maximus L. in the Clyde Sea Area. ICES Journal of Marine Science, 35(3):281–295, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/281/672767.

Corten:1974:RCS

[Cor74] A. Corten. Recent changes in the stock of Celtic Sea herring (Clupea harengus L.). ICES Journal of Marine Science, 35(2): 194–201, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/194/696015.

Campbell:1972:BGC

[CP72] N. A. Campbell and B. F. Phillips. The von Bertalanffy growth curve and its application to capture — recapture data in fisheries biology. ICES Journal of Marine Science, 34(2):295–299, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/295/595536.

Crisp:1971:R

[Cri71] D. J. Crisp. Reviews. ICES Journal of Marine Science, 34 (1):140–141, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/140/602173.

Carlberg:1972:DSA

[CS72] Stig R. Carlberg and C. Bo Skarstedt. Determination of small amounts of non-polar hydrocarbons (oil) in sea water. ICES Journal of Marine Science, 34(3):506–515, October 1972.

CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/506/758557.

Cram:1974:OSS

[CS74] D. L. Cram and F. H. Schülein. Observations on surface-shoaling Cape hake off South West Africa. ICES Journal of Marine Science, 35(3):272–275, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/272/672758.

Chapman:1978:CCC

[CS78] C. J. Chapman and G. L. Smith. Creel catches of crab, Cancer pagurus L. using different baits. ICES Journal of Marine Science, 38(2):226–229, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/226/743209.

Culliney:1974:R

[Cul74] John L. Culliney. Reviews. ICES Journal of Marine Science, 35(3):370–373, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/370/672853.

Curnock:1971:R

[Cur71] W. D. Curnock. Reviews. ICES Journal of Marine Science, 33(3):516–517, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/516/711598.

Currie:1972:R

[Cur72] R. I. Currie. Reviews. ICES Journal of Marine Science, 34 (3):541–542, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/541/758593.

Cushing:1971:DRP

[Cus71a] D. H. Cushing. The dependence of recruitment on parent stock in different groups of fishes. ICES Journal of Marine Science, 33(3):340–362, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/340/711490.

Cushing:1971:Ra

[Cus71b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 34 (1):135–137, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/135/602166.

Cushing:1971:Rb

[Cus71c] D. H. Cushing. Reviews. ICES Journal of Marine Science, 34 (1):142–144, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/142/602176.

Cushing:1972:Ra

[Cus72a] D. H. Cushing. Reviews. ICES Journal of Marine Science, 34 (3):544–545, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/544/758625.

Cushing:1972:Rb

[Cus72b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 34 (3):546–549, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/546/758641.

Cushing:1973:CSE

[Cus73a] D. H. Cushing. Computations with a sonar equation. ICES Journal of Marine Science, 35(1):22–26, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/35/1/22/769684.

Cushing:1973:R

[Cus73b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 35(1):102–103, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/102/769559.

Cushing:1974:RWA

[Cus74a] D. H. Cushing. R. S. Wimpenny (10 August 1898–5 January 1971). ICES Journal of Marine Science, 36(1):1–3, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
1/1/642593.

Cushing:1974:Ra

[Cus74b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 35(2):215, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/215/696062.

Cushing:1974:Rb

[Cus74c] D. H. Cushing. Reviews. ICES Journal of Marine Science, 35(3):376–377, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/376/672882.

Cushing:1974:Rc

[Cus74d] D. H. Cushing. Reviews. ICES Journal of Marine Science, 36(1):95, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/95/642677.

Cushing:1975:NMP

[Cus75a] D. H. Cushing. The natural mortality of the plaice. ICES Journal of Marine Science, 36(2):150–157, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/36/2/150/746314.

Cushing:1975:R

[Cus75b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 36(2):194–195, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/194/746401.

Cushing:1976:PP

[Cus76a] D. H. Cushing. In praise of Petersen. ICES Journal of Marine Science, 36(3):277–281, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/36/3/277/639257.

Cushing:1976:R

[Cus76b] D. H. Cushing. Reviews. ICES Journal of Marine Science, 37(1):114–115, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/1/114/622169.

Cushing:1978:PSA

[Cus78] D. H. Cushing. The present state of acoustic survey. ICES Journal of Marine Science, 38(1):28–32, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/1/28/630218.

Dare:1974:DCM

[Dar74] P. J. Dare. Damage caused to mussels (Mytilus edulis L.) by dredging and mechanized sorting. ICES Journal of Marine Science, 35(3):296–299, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/296/672769.

Dare:1977:EEP

[Dar77] P. J. Dare. Enzyme extraction of the parasitic copepod Mytilicola intestinalis Steuer from mussels, Mytilus edulis L. ICES Journal of Marine Science, 37(2):170–172, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/37/2/170/603444.

Davies:1974:MMS

[Dav74] Gwion Davies. A method for monitoring the spatfall of mussels (Mytilus edulis L.). ICES Journal of Marine Science, 36 (1):27–34, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/27/642600.

DeAngelis:1979:GSR

[DC79] D. L. DeAngelis and S. W. Christensen. A general stockrecruitment curve. ICES Journal of Marine Science, 38(3):324–325, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/3/324/646615.

deCiechomski:1971:ITN

[dCdV71] Janina Dz. de Ciechomski and Weiss G. de Vigo. The influence of the temperature on the number of vertebrae in the Argentine anchovy, Mengraulis anchoita (Hubbs and Martni). ICES Journal of Marine Science, 34(1):37–42, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/1/37/602212.

Rosado:1971:LTC

[dCR71] J. M. de Campos Rosado. Long-term changes in abundance of yellowfin and skipjack off the Coast of Angola. ICES Journal of Marine Science, 34(1):65–75, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/1/65/602229.

Rosado:1973:CEA

[dCR73] J. M. de Campos Rosado. The catch and effort of the Angolan bait-boat tuna fishery and its influence in reducing the supporting yellowfin population. ICES Journal of Marine Science, 35(1):52–60, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/52/769762.

DeCamposRosado:1974:EFS

[De 74] J. M. De Campos Rosado. Evaluating fish stock changes from the fragmentary catch and effort data of the Angolan fisheries. ICES Journal of Marine Science, 35(2):180–184, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/180/696000.

Dare:1976:ESG

[DE76] P. J. Dare and D. B. Edwards. Experiments on the survival, growth and yield of relaid seed mussels (Mytilus edulis L.) in the Menai Straits, North Wales. ICES Journal of Marine Science, 37(1):16–28, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/1/16/622175.

Dickson:1972:RBO

[Dic72] R. R. Dickson. On the relationship between ocean transparency and the depth of sonic scattering layers in the North Atlantic. ICES Journal of Marine Science, 34(3):416–422, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/416/758524.

Dickson:1976:FTS

[Dic76] Robert R. Dickson. Field tests of seabed drifters, using sectorscanning sonar techniques. ICES Journal of Marine Science, 37(1):

3–15, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/37/1/3/622181.

Dickson:1978:EDC

[DLMS78] R. R. Dickson, D. N. Langhorne, R. S. Millner, and E. G. Shreeve. An examination of a dredged channel using sector scanning sonar in side-scan mode. ICES Journal of Marine Science, 38(1):41–47, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/41/630231.

Dooley:1974:HCC

[Doo74] H. D. Dooley. Hypotheses concerning the circulation of the northern North Sea. ICES Journal of Marine Science, 36(1):54–61, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/1/54/642629.

Dow:1972:FGM

[Dow72] Robert L. Dow. Fluctuations in Gulf of Maine Sea temperature and specific molluscan abundance. ICES Journal of Marine Science, 34 (3):532–534, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/532/758575.

Dow:1977:ECC

[Dow77a] Robert L. Dow. Effects of climatic cycles on the relative abundance and availability of commercial marine and estuarine species. ICES Journal of Marine Science, 37(3):274–280, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
3/274/637233.

Dow:1977:RSS

[Dow77b] Robert L. Dow. Relationship of sea surface temperature to American and European lobster landings. ICES Journal of Marine Science, 37(2):186–191, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/37/2/186/603448.

Dow:1978:ESS

[Dow78] Robert L. Dow. Effects of sea-surface temperature cycles on landings of American, European and Norway lobsters. ICES Journal of Marine Science, 38(2):271–272, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/2/271/743264.

DuBuit:1975:ERT

[Du 75] M. H. Du Buit. Étude de la relation taille/poids chez Raja naevus (Rajidae). coefficient de condition. (French) [Study of the size/weight relation in Raja neevus (Rajidae). Condition coefficient]. ICES Journal of Marine Science, 36(2):166–169, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/166/746340.

DuBuit:1977:ACR

[Du 77] Marie Henriette Du Buit. Âge et croissance de Raja batis et de Raja naevus en Mer Celtique. (French) [Age and growth of Raja batis and Raja neevus in Celtic Sea]. ICES Journal of Marine Science, 37(3):261–265, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/3/261/637222.

deVeen:1970:OPP

[dV70] J. F. de Veen. On the orientation of the plaice (Pleuronectes platessa L.) I. Evidence for orientating factors derived from the ices transplantation experiments in the years 1904–1909. ICES Journal of Marine Science, 33(2):192–227, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/2/192/800680.

deVeen:1976:CSB

[dV76] J. F. de Veen. On changes in some biological parameters in the North Sea sole (Solea solea L.). ICES Journal of Marine Science, 37(1):60–90, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/1/60/622191.

Davies:1977:FSA

[DVB77] Isaac E. Davies, Robert J. Vent, and John C. Brown. Fish school acoustic target strength. ICES Journal of Marine Science, 37(3): 288–292, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/288/637247.

Ellett:1972:R

[Ell72] D. J. Ellett. Reviews. ICES Journal of Marine Science, 34 (2):313–315, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/313/595542.

Engen:1976:NES

[Eng76] Steinar Engen. A note on the estimation of the species-area curve. ICES Journal of Marine Science, 36(3):286–288, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/3/286/639267.

El-Zarka:1970:SWB

[EZKS70] S. El-Zarka, R. Koura, and A. H. Shaheen. Selectivity of wire basket traps for tilapia (T. nilotica, T. galilaea and T. zillii). ICES Journal of Marine Science, 33(2):282–291, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/282/800710.

Fahy:1972:ITC

[Fah72] William E. Fahy. Influence of temperature change on number of vertebrae and caudal fin rays in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 34(2):217–231, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/217/595522.

Fahy:1976:MTF

[Fah76] W. E. Fahy. The morphological time of fixation of the total number of vertebrae in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 36(3):243–250, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/36/3/243/639239.

Fahy:1978:ICU

[Fah78] W. E. Fahy. The influence of crowding upon the total number of vertebrae developing in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 38(2):252–256, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/2/252/743241.

Fahy:1979:ITC

[Fah79] W. E. Fahy. The influence of temperature change on number of anal fin rays developing in Fundulus majalis (Walbaum). ICES Journal of Marine Science, 38(3):280–285, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/3/280/646579.

Flemming:1979:R

[Fle79] N. C. Flemming. Reviews. ICES Journal of Marine Science, 38 (3):346–347, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/346/646632.

Fahy:1977:DSI

[FO77] W. E. Fahy and R. K. O’Hara. Does salinity influence the number of vertebrae developing in fishes? ICES Journal of Marine Science, 37(2):156–161, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/156/603438.

Forster:1972:SMB

[For72] J. R. M. Forster. Some methods of binding prawn diets and their effects on growth and assimilation. ICES Journal of Marine Science, 34(2):200–216, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/200/595519.

Fraser:1970:ECP

[Fra70] J. H. Fraser. The ecology of the ctenophore Pleurobrachia pileus in Scottish waters. ICES Journal of Marine Science, 33(2):149–168, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
33/2/149/800659.

Fraser:1971:R

[Fra71] J. H. Fraser. Reviews. ICES Journal of Marine Science, 34 (1):147–149, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/147/602185.

Frost:1978:SCS

[Fro78] Winifred E. Frost. The scales of the charr, (Salvelinus willughbii Günther), in Windermere, and their use for the determination of age and growth. ICES Journal of Marine Science, 38 (2):208–215, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/208/743181.

Garrod:1976:R

[Gar76] D. J. Garrod. Reviews. ICES Journal of Marine Science, 36 (3):297, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/297/639280.

Garrod:1977:R

[Gar77] D. J. Garrod. Reviews. ICES Journal of Marine Science, 37 (2):195, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/195/603456.

Gibson:1978:RCP

[Gib78] R. John Gibson. Recent changes in the population of juvenile Atlantic salmon in the Matamek River, Quebec, Canada. ICES Journal of Marine Science, 38(2):201–207, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
2/201/743174.

Grasshoff:1972:NSD

[GJ72] K. Grasshoff and H. Johannsen. A new sensitive and direct method for the automatic determination of ammonia in sea water. ICES Journal of Marine Science, 34(3):516–521, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/516/758562.

Garrod:1974:SRR

[GJ74a] D. J. Garrod and B. W. Jones. Stock and recruitment relationship in the Northeast Arctic cod stock and the implications for management of the stock. ICES Journal of Marine Science, 36 (1):35–41, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/35/642603.

Grasshoff:1974:NTM

[GJ74b] K. Grasshoff and H. Johannsen. Notes on techniques and methods for sea water analysis: a critical review of the method by benesch and mangelsdorf for the colorimetric determination of ammonia in sea water. ICES Journal of Marine Science, 36(1):90–92, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/1/90/642666.

Gee:1977:RBI

[GMD77] J. M. Gee, Linda Maddock, and J. T. Davey. The relationship between infestation by Mytilicola intestinalis, Steuer (Copepoda, Cyclopoidea) and the condition index of Mytilus edulis in southwest England. ICES Journal of Marine Science, 37(3):300–308, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
37/3/300/637269.

Gould:1974:R

[Gou74] W. J. Gould. Reviews. ICES Journal of Marine Science, 35 (2):214–215, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/214/696058.

Gardner:1974:MAA

[GR74] Doris Gardner and J. P. Riley. Mercury in the Atlantic around Iceland. ICES Journal of Marine Science, 35(2):202–204, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/2/202/696018.

Grainger:1978:HAW

[Gra78] R. J. R. Grainger. Herring abundance off the west of Ireland in relation to oceanographic variation. ICES Journal of Marine Science, 38(2):180–188, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/180/743153.

Gruffydd:1974:ENM

[Gru74a] Ll. D. Gruffydd. An estimate of natural mortality in an unfished population of the scallop Pecten maximus (L.). ICES Journal of Marine Science, 35(2):209–210, February 1974. CODEN ICESEC.

ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/35/2/209/696029.

Gruffydd:1974:ICE

[Gru74b] Ll. D. Gruffydd. The influence of certain environmental factors on the maximum length of the scallop, Pecten maximus L. ICES Journal of Marine Science, 35(3):300–302, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/300/672772.

Gabbott:1974:NRB

[GS74] P. A. Gabbott and R. R. Stephenson. A note on the relationship between the dry weight condition index and the glycogen content of adult oysters (Ostrea edulis L.) kept in the laboratory. ICES Journal of Marine Science, 35(3):359–361, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/359/672829.

Gueredrat:1970:NMN

[Gue70] J.-A. Gueredrat. Note sur les migrations nycthémérales de Labidocera acutifrons (Crustacé — Copépode). (French) [Note on the nycthemeral migrations of Labidocera acutifrons (Crustacea — Copepod)]. ICES Journal of Marine Science, 33(2):293–296, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/293/800725.

Gulland:1971:R

[Gul71a] J. A. Gulland. Reviews. ICES Journal of Marine Science, 33 (3):518–519, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/518/711606.

Gulland:1971:SFM

[Gul71b] J. A. Gulland. Science and fishery management. ICES Journal of Marine Science, 33(3):471–477, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/3/471/711535.

Gulland:1972:R

[Gul72] J. A. Gulland. Reviews. ICES Journal of Marine Science, 34 (3):543–544, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/543/758603.

Gulland:1977:SFS

[Gul77] J. A. Gulland. The stability of fish stocks. ICES Journal of Marine Science, 37(3):199–204, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/3/199/637193.

Gulland:1979:R

[Gul79] J. A. Gulland. Reviews. ICES Journal of Marine Science, 38 (3):347–348, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/347/646635.

Greer-Walker:1970:GDS

[GW70] M. Greer-Walker. Growth and development of the skeletal muscle fibres of the cod (Gadus morhua L.). ICES Journal of Marine Science, 33(2):228–244, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/2/228/800686.

Gabbott:1971:CCI

[GW71] P. A. Gabbott and A. J. M. Walker. Changes in the condition index and biochemical content of adult oysters (Ostreaedulis L.) maintained under hatchery conditions. ICES Journal of Marine Science, 34(1):98–105, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/1/98/602239.

Grasshoff:1972:MVM

[GW72] K. Grasshoff and A. Wenck. A modern version of the Mohr– Knudsen titration for the chlorinity of sea water. ICES Journal of Marine Science, 34(3):522–528, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/3/522/758566.

Harding:1971:FEH

[HA71] D. Harding and G. P. Arnold. Flume experiments on the hydrodynamics of the Lowestoft high-speed plankton samplers: I. ICES Journal of Marine Science, 34(1):24–36, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/34/1/24/602207.

Hampson:1977:RBT

[Ham77] B. L. Hampson. Relationship between total ammonia and free ammonia in terrestrial and ocean waters. ICES Journal of Marine Science, 37(2):117–122, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/2/117/603418.

Hancock:1974:AAM

[Han74] D. A. Hancock. Attraction and avoidance in marine invertebrates — their possible role in developing an artificial bait. ICES Journal of Marine Science, 35(3):328–331, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/35/3/328/672814.

Hartmann:1970:JSP

[Har70] Jürgen Hartmann. Juvenile saury pike (Scomberesox saurus Walb.), an example of ichthyoneuston. ICES Journal of Marine Science, 33(2):245–255, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/245/800695.

Harris:1975:EDD

[Har75] J. G. K. Harris. The effect of density-dependent mortality on the shape of the stock and recruitment curve. ICES Journal of Marine Science, 36(2):144–149, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/36/2/144/746304.

Hargreaves:1976:ETN

[Har76] P. M. Hargreaves. Echo-traces from the North-eastern Atlantic. ICES Journal of Marine Science, 37(1):46–59, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/37/1/46/622188.

Hawkins:1978:R

[Haw78] A. D. Hawkins. Reviews. ICES Journal of Marine Science, 38(1):125–126, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/125/630206.

Horwood:1976:NTS

[HD76] J. W. Horwood and R. M. Driver. A note on a theoretical subsampling distribution of macroplankton. ICES Journal of Marine Science, 36(3):274–276, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/36/3/274/639251.

Head:1975:PCF

[Hea75] Anna Head. Preparation of copper ferrocyanide for extraction from sea water: a warning. ICES Journal of Marine Science, 36(2):193, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/2/193/746394.

Hemmings:1971:R

[Hem71] C. C. Hemmings. Reviews. ICES Journal of Marine Science, 34 (1):141–142, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/141/602175.

Hemmings:1972:UPF

[Hem72] C. C. Hemmings. Underwater photography in fisheries research. ICES Journal of Marine Science, 34(3):466–484, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/466/758540.

Hennig:1974:ELA

[Hen74] H. F. K. O. Hennig. The effect of a larval Anisakis (Nematoda: Ascaroidea) on the South West African anchovy, Engraulis capensis. ICES Journal of Marine Science, 35(2):185–188, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/185/696005.

Hepper:1972:GML

[Hep72] B. T. Hepper. The growth at moulting of lobsters, Homarus vulgaris Milne-edwards, in the Menai Straits, North Wales. ICES Journal of Marine Science, 34(2):169–173, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/169/595514.

Herring:1974:R

[Her74] Peter J. Herring. Reviews. ICES Journal of Marine Science, 35(3):375–376, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/375/672876.

Hepper:1978:FRE

[HG78] B. T. Hepper and C. J. Gough. Fecundity and rate of embryonic development of the lobster, Homarus gammarus (L), off the coast of North Wales. ICES Journal of Marine Science, 38(1):54–57, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/54/630241.

Hecht:1971:NSP

[HH71a] A. Hecht and P. Hughes. A note on short period temperature fluctuations in the North Sea. ICES Journal of Marine Science, 34(1):6–17, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/6/602225.

Hislop:1971:ODS

[HH71b] J. R. G. Hislop and C. C. Hemmings. Observations by divers on the survival of tagged and untagged haddock Melanogrammus aeglefinus (L.) after capture by trawl or Danish seine net. ICES Journal of Marine Science, 33(3):428–437, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/3/428/711521.

Hill:1974:CSS

[HH74a] H. W. Hill and J. W. Horwood. A computer simulation of surface drifter returns. ICES Journal of Marine Science, 35(2): 158–164, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/158/695984.

Hislop:1974:FWM

[HH74b] J. R. G. Hislop and W. B. Hall. The fecundity of whiting, Merlangius merlangus (L.) in the North Sea, the minch and at Iceland. ICES Journal of Marine Science, 36(1):42–49, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
1/42/642613.

Holland:1976:GCW

[HH76a] D. L. Holland and P. J. Hannant. The glycogen content in winter and summer of oysters, Ostrea edulis L., of different ages. ICES Journal of Marine Science, 36(3):240–242, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
3/240/639234.

Houghton:1976:PEC

[HH76b] R. G. Houghton and D. Harding. The plaice of the English Channel: spawning and migration. ICES Journal of Marine Science, 36 (3):229–239, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/229/639232.

Holden:1979:MTG

[HH79] M. J. Holden and R. G. Horrod. The migrations of tope, Galeorhinus galeus (L), in the eastern North Atlantic as determined by tagging. ICES Journal of Marine Science, 38(3):314–317, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/3/314/646601.

Hillden:1978:ALK

[Hil78] Nils-Ove Hilldén. An age–length key for Ctenolabrus rupestris (L) in Swedish waters. ICES Journal of Marine Science, 38(2): 270–271, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/270/743259.

Hislop:1975:BGW

[His75] J. R. G. Hislop. The breeding and growth of whiting, Merlangius merlangus in captivity. ICES Journal of Marine Science, 36(2):119–127, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/119/746277.

Hislop:1976:E

[His76] J. R. G. Hislop. Erratum. ICES Journal of Marine Science, 36 (3):301–302, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/301/639299.

Hojerslev:1977:COI

[HJK77] N. Højerslev, N. G. Jerlov, and G. Kullenberg. Colour of the ocean as an indicator in photosynthetic studies. ICES Journal of Marine Science, 37(3):313–316, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/3/313/637296.

Hislop:1976:PSW

[HM76] J. R. G. Hislop and K. MacKenzie. Population studies of the whiting Merlangius merlangus (L.) of the northern North Sea. ICES Journal of Marine Science, 37(1):98–110, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/37/1/98/622200.

Hojerslv:1978:DMA

[Høj78] Niels K. Højerslv. Daylight measurements appropriate for photosynthetic studies in natural sea waters. ICES Journal of Marine Science, 38(2):131–146, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/38/2/131/743127.

Holden:1971:R

[Hol71] M. J. Holden. Reviews. ICES Journal of Marine Science, 34 (1):145–146, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/145/602181.

Holden:1972:GRR

[Hol72] M. J. Holden. The growth rates of Raja brachyura, R. clavata and R. montagui as determined from tagging data. ICES Journal of Marine Science, 34(2):161–168, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/2/161/595512.

Holden:1975:FRC

[Hol75a] M. J. Holden. The fecundity of Raja clavata in British waters. ICES Journal of Marine Science, 36(2):110–118, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/110/746272.

Holthe:1975:MCO

[Hol75b] Torleif Holthe. A method for the calculation of ordinate values of the cumulative species–area curve. ICES Journal of Marine Science, 36(2):183–184, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/183/746368.

Horwood:1975:R

[Hor75] J. W. Horwood. Reviews. ICES Journal of Marine Science, 36(2):195–196, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/195/746407.

Houghton:1977:FPT

[Hou77] R. G. Houghton. The fishing power of trawlers in the western English Channel between 1965 and 1968. ICES Journal of Marine Science, 37(2):130–136, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/2/130/603424.

Howell:1973:MFC

[How73] B. R. Howell. Marine fish culture in Britain VIII. A marine rotifer, Brachionus plicatilis Muller, and the larvae of the mussel, Mytilus edulis L., as foods for larval flatfish. ICES Journal of Marine Science, 35(1):1–6, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/1/769534.

Holden:1971:REL

[HRH71] M. J. Holden, D. W. Rout, and C. N. Humphreys. The rate of egg laying by three species of ray. ICES Journal of Marine Science, 33(3):335–339, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/335/711482.

Helm:1972:IRA

[HS72] M. M. Helm and B. E. Spencer. The importance of the rate of aeration in hatchery cultures of the larvae of Ostrea edulis L. ICES Journal of Marine Science, 34(2):244–255, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/244/595528.

Harding:1971:NCM

[HSTM71] D. Harding, E. Shreeve, D. S. Tungate, and D. Mummery. A net-changing mechanism for the Lowestoft multipurpose sampler. ICES Journal of Marine Science, 33(3):483–491, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
3/483/711548.

Holden:1974:FRC

[HT74] M. J. Holden and R. N. Tucker. The food of Raja clavata Linnaeus 1758, Raja montagui Fowler 1910, Raja naevus Müller and Henle 1841 and Raja brachyura Lafont 1873 in British waters. ICES Journal of Marine Science, 35(2):189–193, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/189/696011.

Holden:1973:AVS

[HV73] M. J. Holden and M. R. Vince. Age validation studies on the centra of Raja clavata using tetracycline. ICES Journal of Marine Science, 35(1):13–17, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/35/1/13/769662.

Holland:1975:BCC

[HW75] D. L. Holland and G. Walker. The biochemical composition of the cypris larva of the barnacle Balanus balanoides L. ICES Journal of Marine Science, 36(2):162–165, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/36/2/162/746332.

Ibanez:1972:IDE

[Iba72] F. Ibanez. Interprétation de données Écologiques par l’Analyse des composantes principales: Écologie planctonique de la Mer du Nord. (French) [Interpretation of ecological data by principal component analysis: Plankton ecology of the North Sea]. ICES Journal of Marine Science, 34(3):323–340, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/3/323/758457.

Iles:1971:EAG

[Ile71a] T. D. Iles. Ecological aspects of growth in African cichlid fishes. ICES Journal of Marine Science, 33(3):363–385, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
3/363/711499.

Iles:1971:GSN

[Ile71b] T. D. Iles. Growth studies on North Sea herring III. The growth of East Anglian herring during the adult stage of the life history for the years 1940 to 1967. ICES Journal of Marine Science, 33(3):386–420, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/386/711505.

Jamieson:1971:R

[Jam71] Alan Jamieson. Reviews. ICES Journal of Marine Science, 33(3):519–520, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/519/711612.

Jones:1979:STS

[JAWS79] F. R. Harden Jones, G. P. Arnold, M. Greer Walker, and P. Scholes. Selective tidal stream transport and the migration of plaice (Pleuronectes platessa L.) in the southern North Sea. ICES Journal of Marine Science, 38(3):331–337, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/3/331/646623.

Jerlov:1971:R

[Jer71] N. G. Jerlov. Reviews. ICES Journal of Marine Science, 34 (1):151, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/151/602192.

Jerlov:1975:LPC

[Jer75] N. G. Jerlov. Long period changes of the optical properties of the Baltic. ICES Journal of Marine Science, 36(2):188–190, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/188/746382.

Jerlov:1977:CSW

[Jer77] N. G. Jerlov. Classification of sea water in terms of quanta irradiance. ICES Journal of Marine Science, 37(3):281–287, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
3/281/637239.

Jones:1972:IGH

[JH72] R. Jones and J. R. G. Hislop. Investigations into the growth of haddock, Melanogrammus aeglefinus (L) and whiting, Merlangius merlangus (L) in aquaria. ICES Journal of Marine Science, 34 (2):174–189, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/174/595516.

Jones:1978:FOR

[JH78] R. Jones and J. R. G. Hislop. Further observations on the relation between food intake and growth of gadoids in captivity. ICES Journal of Marine Science, 38(2):244–251, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
2/244/743234.

Jones:1972:IAL

[Jon72] A. J. Jones. An inexpensive apparatus for the large-scale hatching of Artemia salina L. ICES Journal of Marine Science, 34 (3):351–356, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/351/758466.

Jones:1973:TBF

[Jon73] F. R. Harden Jones. Tail beat frequency, amplitude, and swimming speed of a shark tracked by sector scanning sonar. ICES Journal of Marine Science, 35(1):95–97, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/1/95/769856.

Jones:1974:REF

[Jon74] R. Jones. The rate of elimination of food from the stomachs of haddock Melanogrammus aeglefinus, cod Gadus morhua and whiting Merlangius merlangus. ICES Journal of Marine Science, 35(3):225–243, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/225/672742.

Jones:1977:R

[Jon77] F. R. Harden Jones. Reviews. ICES Journal of Marine Science, 37(2):195, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/195/603459.

Jones:1978:EFC

[Jon78] R. Jones. Estimates of the food consumption of haddock (Melanogrammus aeglefinus) and cod (Gadus morhua). ICES Journal of Marine Science, 38(1):18–27, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/1/18/630211.

Jones:1970:PMS

[JS70] Margaret Jones and C. P. Spencer. The phytoplankton of the Menai Straits. ICES Journal of Marine Science, 33(2):169–180, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/169/800668.

Jones:1974:ELT

[JS74a] F. R. Harden Jones and P. Scholes. The effect of low temperature on cod, Gadus morhua. ICES Journal of Marine Science, 35(3):258–271, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/258/672752.

Jones:1974:EDD

[JS74b] F. R. Harden Jones and P. Scholes. The effect of the door-to-door tickler chain on the catch-rate of plaice (Pleuronectes platessa L.) taken by an otter trawl. ICES Journal of Marine Science, 35 (2):210–212, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/210/696032.

Johnson:1975:SFI

[JUH75] Allyn G. Johnson, Fred M. Utter, and Harold O. Hodgins. Study of the feasibility of immunochemical methods for identification of pleuronectid eggs. ICES Journal of Marine Science, 36(2):158–161, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/2/158/746322.

Jones:1973:MWS

[JWA73] F. R. Harden Jones, M. Greer Walker, and G. P. Arnold. The movements of a woodhead seabed drifter tracked by sector scanning sonar. ICES Journal of Marine Science, 35(1):87–92, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
1/87/769822.

Kasahara:1971:R

[Kas71] Hiroshi Kasahara. Reviews. ICES Journal of Marine Science, 34 (1):152–153, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/152/602195.

Kipling:1978:WLR

[KF78] Charlotte Kipling and Winifred E. Frost. The weight–length relationship of charr, (Salvelinus willughbii Günther), in Windermere. ICES Journal of Marine Science, 38(2):216–219, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/2/216/743192.

Kimura:1978:LME

[Kim78] Daniel K. Kimura. Logistic model for estimating selection ogives from catches of codends whose ogives overlap. ICES Journal of Marine Science, 38(1):116–119, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/38/1/116/630178.

Kinne:1973:R

[Kin73] O. Kinne. Reviews. ICES Journal of Marine Science, 35(1):103, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/103/769566.

Kirkerud:1974:MRS

[Kir74] Lars A. Kirkerud. A method for the rapid sorting of plankton into a number of size groups. ICES Journal of Marine Science, 35(3):367–369, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/367/672847.

Korringa:1979:BH

[Kor79] P. Korringa. Berend havinga. ICES Journal of Marine Science, 38(3):277–279, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/277/646575.

Kroger:1974:LFT

[KSL74] R. L. Kroger, G. B. Sekavec, and E. J. Levi. Loss of ferromagnetic tags when injected into the air bladder and abdominal cavity of Atlantic menhaden, Brevoortia tyrannus. ICES Journal of Marine Science, 35(3):244–248, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/35/3/244/672747.

Larkin:1971:R

[Lar71] P. A. Larkin. Reviews. ICES Journal of Marine Science, 34 (1):149–150, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/149/602188.

Last:1972:EDF

[Las72] J. M. Last. Egg development, fecundity and growth of Oikopleura dioica Fol in the North Sea. ICES Journal of Marine Science, 34(2):232–237, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/232/595524.

Last:1979:FLT

[Las79] J. M. Last. The food of larval turbot Scophthalmus maximus L. from the west central North Sea. ICES Journal of Marine Science, 38(3):308–313, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/308/646597.

Laurec:1977:AEP

[Lau77] Alain Laurec. Analyse et estimation des puissances de pêche. ICES Journal of Marine Science, 37(2):173–185, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
2/173/603446.

Lockwood:1975:FOE

[LdBD75] Stephen J. Lockwood and C. de B. Daly. Further observations on the effects of preservation in 4% neutral formalin on the length and weight of 0-group flatfish. ICES Journal of Marine Science, 36(2):170–175, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/170/746349.

Lebedeva:1972:MLD

[Leb72] L. P. Lebedeva. A model of the latitudinal distribution of the number of species of phytoplankton in the sea. ICES Journal of Marine Science, 34(3):341–350, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/3/341/758461.

Lee:1974:GDN

[Lee74] A. J. Lee. Günter Dietrich: 15 November 1911–2 October 1972. ICES Journal of Marine Science, 36(1):4–6, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
1/4/642608.

Lenz:1973:OCB

[Len73] Jurgen Lenz. On the occurrence of the ctenophore Bolinopsis infundibulum (O. F. Müller) in the Western Baltic. ICES Journal of Marine Science, 35(1):32–35, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/1/32/769721.

Letaconnoux:1970:ED

[Let70] R. Letaconnoux. Edouard Le Danois (1887–1968). ICES Journal of Marine Science, 33(2):122–123, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/2/122/800620.

Letaconnoux:1971:R

[Let71] R. Letaconnoux. Reviews. ICES Journal of Marine Science, 34 (1):156–157, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/156/602199.

Lindquist:1974:MMN

[LH74] Armin Lindquist and Lennart Hannerz. Migrations of the mackerel in the northern North Sea and in the Skagerak. ICES Journal of Marine Science, 35(3):276–280, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/3/276/672764.

Lindenberg:1973:MSA

[Lin73] John G. Lindenberg. Macroplankton sorting aid. ICES Journal of Marine Science, 35(1):97–98, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/1/97/769866.

Liss:1974:R

[Lis74] P. S. Liss. Reviews. ICES Journal of Marine Science, 36(1): 97–98, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/97/642694.

Liss:1976:R

[Lis76] P. S. Liss. Reviews. ICES Journal of Marine Science, 37 (1):112–114, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/1/112/622165.

Lloyd:1971:PPC

[Llo71] I. J. Lloyd. Primary production off the Coast of North-West Africa. ICES Journal of Marine Science, 33(3):312–323, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
3/312/711468.

Langton:1974:ECV

[LM74] R. W. Langton and G. U. McKay. The effect of continuous versus discontinuous feeding on the growth of hatchery reared spat of Crassostrea gigas Thunberg. ICES Journal of Marine Science, 35(3):361–363, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/361/672833.

LaBar:1978:SMH

[LMF78] George W. LaBar, James D. McCleave, and Stephen M. Fried. Seaward migration of hatchery-reared Atlantic salmon (Salmo salar) smolts in the Penobscot River estuary, Maine: open-water movements. ICES Journal of Marine Science, 38(2):257–269, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
2/257/743248.

Lockyer:1972:ASM

[Loc72] C. Lockyer. The age at sexual maturity of the southern fin whale (Balaenoptera physalus) using annual layer counts in the ear plug. ICES Journal of Marine Science, 34(2):276–294, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/276/595535.

Lockwood:1973:WLC

[Loc73] S. J. Lockwood. Weight and length changes of 0-group plaice (Pleuronectes platessa L.) after preservation in 4% neutral formalin. ICES Journal of Marine Science, 35(1):100–101, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
1/100/769551.

Lockwood:1974:UMG

[Loc74a] Stephen J. Lockwood. The use of a modified Gulf V plankton sampler from a small open boat. ICES Journal of Marine Science, 35(2):171–174, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/171/695994.

Lockwood:1974:UBG

[Loc74b] Stephen J. Lockwood. The use of the von Bertalanffy growth equation to describe the seasonal growth of fish. ICES Journal of Marine Science, 35(2):175–179, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/2/175/695996.

Lockyer:1974:IEP

[Loc74c] C. Lockyer. Investigation of the ear plug of the southern sei whale, Balaenoptera borealis, as a valid means of determining age. ICES Journal of Marine Science, 36(1):71–81, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/36/1/71/642646.

Lockyer:1976:BWS

[Loc76] C. Lockyer. Body weights of some species of large whales. ICES Journal of Marine Science, 36(3):259–273, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
3/259/639247.

Longhurst:1972:R

[Lon72] Alan R. Longhurst. Reviews. ICES Journal of Marine Science, 34(3):539–541, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/3/539/758584.

Loncarevic:1974:R

[Lon74a] B. D. Loncarevic. Reviews. ICES Journal of Marine Science, 35 (2):217–220, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/217/696075.

Longhurst:1974:Ra

[Lon74b] A. R. Longhurst. Reviews. ICES Journal of Marine Science, 35(3):377–378, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/377/672885.

Longhurst:1974:Rb

[Lon74c] A. R. Longhurst. Reviews. ICES Journal of Marine Science, 36 (1):95–97, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/95/642683.

Lorenzen:1971:CDS

[Lor71] Carl J. Lorenzen. Continuity in the distribution of surface chlorophyll. ICES Journal of Marine Science, 34(1):18–23, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
1/18/602205.

Lorenzen:1972:ELO

[Lor72] Carl J. Lorenzen. Extinction of light in the ocean by phytoplankton. ICES Journal of Marine Science, 34(2):262–267, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/262/595532.

Love:1974:CSC

[Lov74] R. M. Love. Colour stability in cod (Gadus morhua L.) from different grounds. ICES Journal of Marine Science, 35(2):207–209, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/35/2/207/696025.

Lucas:1970:HW

[LPR70] C. E. Lucas, B. B. Parrish, and B. B. Rae. Henry Wood (1894– 1969). ICES Journal of Marine Science, 33(2):119–121, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/119/800614.

Laurence:1976:ETS

[LR76] Geoffrey C. Laurence and Carolyn A. Rogers. Effects of temperature and salinity on comparative embryo development and mortality of Atlantic cod (Gadus morhua L.) and haddock (Melanogrammus aeglefinus (L.)). ICES Journal of Marine Science, 36(3): 220–228, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/220/639227.

LeCorre:1978:DAE

[LT78] Pierre Le Corre and Paul Tréguer. Dosage de l’ammonium dans l’eau de mer: comparaison entre deux méthodes d’analyse automatique. (French) [Determination of ammonium in seawater: comparison between two automatic analysis methods]. ICES Journal of Marine Science, 38(2):147–153, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/2/147/743136.

Lythgoe:1978:R

[Lyt78] J. N. Lythgoe. Reviews. ICES Journal of Marine Science, 38(1): 124, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-

9289 (electronic). URL http://academic.oup.com/icesjms/
article/38/1/124/630196.

Marchand:1974:CFP

[Mar74] Michel Marchand. Considérations sur les formes physico-chimiques du cobalt, manganèse, zinc, chrome et fer dans une eau de mer enrichie ou non de matière organique. (French) [Considerations on the physicochemical forms of cobalt, manganese, zinc, chromium and iron in seawater enriched or not with organic matter]. ICES Journal of Marine Science, 35(2):130–142, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/130/695974.

Mauchline:1977:GSC

[Mau77] J. Mauchline. Growth of shrimps, crabs and lobsters — an assessment. ICES Journal of Marine Science, 37(2):162–169, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
2/162/603441.

McCave:1974:R

[McC74] I. N. McCave. Reviews. ICES Journal of Marine Science, 35(3):373, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/3/373/672865.

Meshal:1977:PSI

[Mes77] Amin H. Meshal. The problem of the salinity increase in Lake Qarun (Egypt) and a proposed solution. ICES Journal of Marine Science, 37(2):137–143, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/2/137/603426.

Miller:1978:ECP

[Mil78] Robert J. Miller. Entry of Cancer productus to baited traps. ICES Journal of Marine Science, 38(2):220–225, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
2/220/743200.

Miller:1979:SCT

[Mil79] Robert J. Miller. Saturation of crab traps: reduced entry and escapement. ICES Journal of Marine Science, 38(3):338–345, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/3/338/646627.

Moore:1978:HEI

[MLG78] M. N. Moore, D. M. Lowe, and J. M. Gee. Histopathological effects induced in Mytilus edulis by Mytilicola intestinalis and the histochemistry of the copepod intestinal cells. ICES Journal of Marine Science, 38(1):6–11, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/6/630274.

Morcos:1970:CCS

[Mor70] Selim A. Morcos. Chemical composition of seawater and the variation of calcium and alkalinity. ICES Journal of Marine Science, 33(2):126–133, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/126/800629.

Morcos:1973:TIC

[Mor73] Selim A. Morcos. A table for the ionic composition of sea water based on 1967 atomic weights. ICES Journal of Marine Science, 35(1):94–95, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/94/769846.

Maerker:1976:CDP

[MS76] Marilyn Maerker and Karl-Heinz Szekielda. Chlorophyll determination of phytoplankton: a comparison of in vivo fluorescence with spectrophotometric absorption. ICES Journal of Marine Science, 36(3):217–219, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/217/639223.

Messieh:1970:CUO

[MT70] S. N. Messieh and S. N. Tibbo. A critique on the use of otoliths for ageing Gulf of St. Lawrence herring (Clupea harengus L.). ICES Journal of Marine Science, 33(2):181–191, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
2/181/800674.

Munro:1974:MOA

[Mun74] J. L. Munro. The mode of operation of Antillean fish traps and the relationships between ingress, escapement, catch and soak. ICES Journal of Marine Science, 35(3):337–350, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/337/672821.

Muzinic:1977:SBS

[Muž77] R. Mužinić. On the shoaling behaviour of sardines (Sardina pilchardus) in aquaria. ICES Journal of Marine Science, 37(2): 147–155, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/147/603435.

Nair:1974:R

[Nai74] N. Balakrishnan Nair. A reply. ICES Journal of Marine Science, 35(3):373, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/373/672860.

Nordeng:1971:FPP

[NB71] Hans Nordeng and Per Bratland. Feeding of plaice (Pleuronectes platessa L.) and COD (Gadus morhua L.) larvae. ICES Journal of Marine Science, 34(1):51–57, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/1/51/602220.

Nichols:1972:R

[Nic72] David Nichols. Reviews. ICES Journal of Marine Science, 34(2):317, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/317/595547.

Nielsen:1976:JK

[Nie76] E. Steemann Nielsen. Johannes Krey (1912–1975). ICES Journal of Marine Science, 37(1):1–2, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/1/1/622160.

Nichols:1978:OLS

[NL78] J. H. Nichols and P. Lawton. The occurrence of the larval stages of the lobster Homarus gammarus, (Linnaeus, 1758) off the northeast coast of England in 1976. ICES Journal of Marine Science, 38 (2):234–243, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/234/743225.

Natarajan:1972:NAL

[NN72] P. Natarajan and N. Balakrishnan Nair. On the nature of attachment of Lernaeenicus hemirhamphi Kirtisinghe to the host fish Hemirhamphus xanthopterus C & V. ICES Journal of Marine Science, 34(3):535–538, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/34/3/535/758579.

Nival:1973:DAF

[NN73] Paul Nival and Suzanne Nival. Description d’un appareil à fractionner le plancton utilisable à la mer. (French) [Description of a plankton fractionating device that can be used in the sea]. ICES Journal of Marine Science, 35(1):98–99, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/35/1/98/769878.

Nichols:1978:SIT

[NW78] J. H. Nichols and R. J. Wood. Sorting and identification techniques for clupeoid larvae. ICES Journal of Marine Science, 38(1):48–53, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/48/630237.

Omori:1973:HPS

[OKN73] Makoto Omori, Teruo Konagaya, and Kazuo Noya. History and present status of the fishery of Sergestes lucens (Penaeidea, Decapoda, Crustacea) in Suruga Bay, Japan. ICES Journal of Marine Science, 35(1):61–77, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/1/61/769772.

Paasche:1971:SME

[Paa71] E. Paasche. A simple method for establishing bacteria-free cultures of phototactic flagellates. ICES Journal of Marine Science, 33(3):509–511, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/509/711564.

Paulik:1971:R

[Pau71] G. J. Paulik. Reviews. ICES Journal of Marine Science, 34 (1):137–139, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/137/602170.

Pawson:1976:R

[Paw76] M. G. Pawson. Reviews. ICES Journal of Marine Science, 36 (3):298–299, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/298/639289.

Pawson:1977:RCG

[Paw77] M. G. Pawson. The responses of cod Gadus morhua (L) to chemical attractants in moving water. ICES Journal of Marine Science, 37 (3):316–318, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/316/637304.

Peres:1973:R

[Pér73] J. M. Pérès. Reviews. ICES Journal of Marine Science, 35(1):106, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/106/769621.

Petersen:1977:OSE

[Pet77] G. Høpner Petersen. The occurrence of the sand eel (Ammodytes dubius) in 0.1 m^{2} van Veen grab samples taken during winter in Godhavn Havn, Disko Bugt. ICES Journal of Marine Science, 37 (3):309–310, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/309/637277.

Petersen:1979:ADF

[Pet79] G. Høpner Petersen. On the analysis of dark fixation in primary production computations. ICES Journal of Marine Science, 38 (3):326–330, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/3/326/646618.

Purdom:1971:CIP

[PH71] C. E. Purdom and A. E. Howard. Ciliate infestations: a problem in marine fish farming. ICES Journal of Marine Science, 33(3):511–514, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/511/711574.

Pocklington:1974:RSS

[PH74] R. Pocklington and W. R. Hardstaff. A rapid semiquantitative screening procedure for lignin in marine sediments. ICES Journal of Marine Science, 36(1):92–94, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/36/1/92/642672.

Page-Jones:1971:SIU

[PJ71] R. M. Page-Jones. A salinity indicator for use in marine fish farming. ICES Journal of Marine Science, 34(1):121–123, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
34/1/121/602156.

Platt:1971:APP

[Pla71] Trevor Platt. The annual production by phytoplankton in St. Margaret’s Bay, Nova Scotia. ICES Journal of Marine Science, 33(3):324–333, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/324/711476.

Platt:1976:CPB

[Pla76] N. E. Platt. Codworm — a possible biological indicator of the degree of mixing of Greenland and Iceland cod stocks. ICES Journal of Marine Science, 37(1):41–45, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/1/41/622185.

Portmann:1977:R

[Por77] J. E. Portmann. Reviews. ICES Journal of Marine Science, 37 (3):321–322, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/321/637317.

Portmann:1978:R

[Por78] J. E. Portmann. Reviews. ICES Journal of Marine Science, 38 (2):273–274, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/273/743282.

Postuma:1972:AMS

[Pos72] K. H. Postuma. On the abundance of mackerel (Scomber scombrus L.) in the Northern and North-Eastern North Sea in the period 1959–1969. ICES Journal of Marine Science, 34(3):455–465, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
34/3/455/758535.

Postuma:1974:NHO

[Pos74] K. H. Postuma. The nucleus of the herring otolith as a racial character. ICES Journal of Marine Science, 35(2):121–129, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/121/695970.

Paul:1978:PDG

[PPN78] A. J. Paul, J. M. Paul, and R. A. Nevé. Phytoplankton densities and growth of Mytilus edulis in an Alaskan artificial upwelling system. ICES Journal of Marine Science, 38(1):100–104, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
1/100/630171.

Pollock:1977:GMC

[PR77] D. E. Pollock and M. J. Roscoe. The growth at moulting of crawfish Jasus tristani at Tristan da Cunha, South Atlantic. ICES Journal of Marine Science, 37(2):144–146, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
2/144/603432.

Preston:1973:R

[Pre73] A. Preston. Reviews. ICES Journal of Marine Science, 35(1):105, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/105/769609.

Preston:1974:R

[Pre74] A. Preston. Reviews. ICES Journal of Marine Science, 35 (3):374–375, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/374/672872.

Preston:1977:Ra

[Pre77a] A. Preston. Reviews. ICES Journal of Marine Science, 37(2): 194–195, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/194/603453.

Preston:1977:Rb

[Pre77b] Alan Preston. Reviews. ICES Journal of Marine Science, 37 (3):320, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/320/637313.

Preston:1978:R

[Pre78] A. Preston. Reviews. ICES Journal of Marine Science, 38 (1):124–125, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/124/630201.

Postma:1976:IOT

[PSLC76] H. Postma, A. Svansson, H. Lacombe, and K. Crasshoff. The international oceanographic tables for the solubility of oxygen in sea water. ICES Journal of Marine Science, 36(3):295–296, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/3/295/639275.

Purdom:1974:R

[Pur74] C. E. Purdom. Reviews. ICES Journal of Marine Science, 35 (2):213–214, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/213/696035.

Ramster:1975:DJC

[Ram75] John Ramster. Dr. J. N. Carruthers: 24 November 1895–8 March 1973. ICES Journal of Marine Science, 36(2):101–105, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/101/746255.

Reay:1972:SPO

[Rea72] P. J. Reay. The seasonal pattern of otolith growth and its application to back-calculation studies in Ammodytes tobianus L.

ICES Journal of Marine Science, 34(3):485–504, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/485/758546.

Repelin:1970:ECF

[RG70] R. Repelin and J.-A. Gueredrat. Efficiences comparées de filets à plancton coniques de mêmes dimensions et de mailles différentes. II. Sélectivité de la maille N^{o}000 pour les amphipodes et les copépodes. (French) [Comparative efficiencies of conical plankton nets of same dimensions and different meshes. II. Mesh selectivity N ^{o}000 for amphipods and copepods]. ICES Journal of Marine Science, 33 (2):256–281, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/2/256/800703.

Richard:1971:DRD

[Ric71] Joseph D. Richard. Delayed release device for use in trap fisheries. ICES Journal of Marine Science, 33(3):492–505, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/33/
3/492/711554.

Rommel:1973:POE

[RM73] Sentiel A. Rommel, Jr. and James D. Mccleave. Prediction of oceanic electric fields in relation to fish migration. ICES Journal of Marine Science, 35(1):27–31, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/35/1/27/769694.

Ramster:1978:NTS

[RM78] J. W. Ramster and K. J. Medler. Notes on the tidal streams off orfordness in relation to fish tracking. ICES Journal of Marine Science, 38(1):87–91, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/38/1/87/630294.

Rollefsen:1970:PB

[Rol70] G. Rollefsen. Paul Bjerkan (1874–1968). ICES Journal of Marine Science, 33(2):124–125, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/33/2/124/800623.

Riley:1972:WSD

[RR72] J. D. Riley and J. W. Ramster. Woodhead seabed drifter recoveries and the influence of human, tidal and wind factors. ICES Journal of Marine Science, 34(3):389–415, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/34/3/389/758481.

Ruffman:1974:R

[Ruf74] Alan Ruffman. Reviews. ICES Journal of Marine Science, 35 (2):220–223, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/220/696078.

Ruud:1971:EL

[Ruu71] Johan T. Ruud. Einar Lea 1887–1969. ICES Journal of Marine Science, 33(3):303–307, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/33/3/303/711460.

Roger:1974:SPA

[RW74] S. Roger and T. Wilson. A simple and precise analytical method for the determination of caesium-137 in sea water. ICES Journal of Marine Science, 36(1):87–89, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/36/1/87/642660.

Revie:1974:TTD

[RWHW74] J. Revie, G. Wearden, P. D. Hocking, and D. E. Weston. A twenty-three day twenty-mile echo record of fish behaviour. ICES Journal of Marine Science, 36(1):82–86, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/36/1/82/642652.

Saetersdal:1977:GR

[Sæt77] Gunnar Sætersdal. Gunnar rollefsen. ICES Journal of Marine Science, 37(3):197–198, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/3/197/637187.

Sahlin:1971:NR

[Sah71] Sven Sahlin. Nils Rosén (1882–1970). ICES Journal of Marine Science, 34(1):3–5, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/34/1/3/602210.

Sanders:1977:EMC

[San77] M. J. Sanders. Estimation of mortality coefficients for exploited fish populations in which the sexes exhibit divergent growth rates. ICES Journal of Marine Science, 37(2):192–193, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/37/
2/192/603451.

Saraiva:1973:ERS

[Sar73] Maria C. Saraiva. Étude sur la radio sensibilité d’un Copépode benthique, Tigriopus fulvus (Fischer). ICES Journal of Marine Science, 35(1):7–12, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/7/769781.

Sawyer:1972:TRF

[Saw72] L. J. E. Sawyer. A technique for recording the filtering activity of marine invertebrates. ICES Journal of Marine Science, 34(2):308–312, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/308/595541.

Seaton:1971:IDE

[SB71] D. D. Seaton and R. S. Bailey. The identification and development of the eggs and larvae of the blue whiting Micromesistius poutassou (Risso). ICES Journal of Marine Science, 34(1):76–83, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
34/1/76/602231.

Stewart:1974:SUD

[SC74] Peter A. M. Stewart and George M. Cameron. The safe use by divers of a high current pulse generator in studies of the behaviour of marine fish in electric fields. ICES Journal of Marine Science, 36(1):62–70, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/62/642635.

Schneider:1972:R

[Sch72] Hans Schneider. Reviews. ICES Journal of Marine Science, 34 (2):315–316, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/2/315/595545.

Scott:1973:R

[Sco73] W. B. Scott. Reviews. ICES Journal of Marine Science, 35 (1):103–104, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/1/103/769577.

Szekielda:1976:PRS

[SD76] Karl-Heinz Szekielda and James T. Duvall. Pattern recognition of suspended material. ICES Journal of Marine Science, 36(3): 205–216, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/205/639218.

Schuyf:1971:ILD

[SdG71] A. Schuyf and S. J. de Groot. An inductive locomotion detector for use in diurnal activity experiments in fish. ICES Journal of Marine Science, 34(1):126–131, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/1/126/602160.

Sundnes:1972:SGP

[SG72] Gunnar Sundnes and Trygve Gytre. Swimbladder gas pressure of cod in relation to hydrostatic pressure. ICES Journal of Marine Science, 34(3):529–532, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/3/529/758571.

Shepherd:1979:SEP

[SH79] J. G. Shepherd and J. W. Horwood. The sensitivity of exploited populations to environmental “noise”, and the implications for management. ICES Journal of Marine Science, 38(3):318–323, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/3/318/646608.

Sigel:1971:R

[Sig71] M. Michael Sigel. Reviews. ICES Journal of Marine Science, 34 (1):153–156, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/153/602197.

Simpson:1971:DSB

[Sim71] A. C. Simpson. Diel spawning behaviour in populations of plaice, dab, sprat and pilchard. ICES Journal of Marine Science, 34 (1):58–64, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/34/1/58/602223.

Sissenwine:1974:VRE

[Sis74] Michael P. Sissenwine. Variability in recruitment and equilibrium catch of the Southern New England yellowtail flounder fishery. ICES Journal of Marine Science, 36(1):15–26, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
1/15/642596.

Sanders:1976:FPF

[SM76] M. J. Sanders and A. J. Morgan. Fishing power, fishing effort, density, fishing intensity and fishing mortality. ICES Journal of Marine Science, 37(1):36–40, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/1/36/622183.

Smed:1971:R

[Sme71] Jens Smed. Reviews. ICES Journal of Marine Science, 33 (3):520–521, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/520/711619.

Smed:1978:FH

[Sme78] Jens Smed. Frede Hermann. ICES Journal of Marine Science, 38(1):3–5, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/3/630221.

Smith:1972:ODE

[Smi72] John W. Smith. The occurrence of Diclidophora esmarkii (Monogenea) on Norway pout, Trisopterus esmarkii (Nilsson, 1855), in the Northern North Sea and to the North and West of Scotland. ICES Journal of Marine Science, 34(2):256–261, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/256/595530.

Smith:1976:R

[Smi76] Paul E. Smith. Reviews. ICES Journal of Marine Science, 36 (3):297–298, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/297/639284.

Smith:1978:PSM

[Smi78] Paul E. Smith. Precision of sonar mapping for pelagic fish assessment in the California current. ICES Journal of Marine Science, 38(1):33–40, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/33/630224.

Sorokin:1971:R

[Sor71] Yu. Sorokin. Reviews. ICES Journal of Marine Science, 33 (3):515–516, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/515/711588.

Sorokin:1972:BPP

[Sor72] Yu. I. Sorokin. The bacterial population and the processes of hydrogen sulphide oxidation in the Black Sea. ICES Journal of Marine Science, 34(3):423–454, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/3/423/758528.

Symonds:1971:SSN

[SS71] D. J. Symonds and A. C. Simpson. The survival of small Nephrops returned to the sea during commercial fishing. ICES Journal of Marine Science, 34(1):89–97, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/1/89/602236.

Sundnes:1975:SPS

[SS75] Gunnar Sundnes and Olav Sand. Studies of a physostome swimbladder by resonance frequency analyses. ICES Journal of Marine Science, 36(2):176–182, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/36/2/176/746357.

Szekielda:1974:VSS

[SSS74] K. H. Szekielda, W. E. Shenk, and V. V. Salomonson. Variability of sea-surface temperatures of the southern Indian Ocean. ICES Journal of Marine Science, 35(2):143–148, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
2/143/695978.

Szekielda:1977:SIU

[SST77] Karl-Heinz Szekielda, Dennis J. Suszkowski, and Paul S. Tabor. Skylab investigation of the upwelling off the northwest coast of Africa. ICES Journal of Marine Science, 37(3):205–213, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
37/3/205/637196.

Stark:1971:CPE

[Sta71] A. E. Stark. A computer programme to estimate fishing power by the method of fitting constants. ICES Journal of Marine Science, 33(3):478–482, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/478/711541.

Steedman:1974:LMS

[Ste74a] H. F. Steedman. Laboratory methods in the study of marine zooplankton: a summary report on the results of Joint Working Group 23 of the Scientific Committee on Oceanic Research and the United Nations Educational, Scientific and Cultural Organisation 1968–1972. ICES Journal of Marine Science, 35(3):351–358, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/3/351/672824.

Steele:1974:R

[Ste74b] J. H. Steele. Reviews. ICES Journal of Marine Science, 35 (2):214, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/214/696038.

Stewart:1974:IEE

[Ste74c] Peter A. M. Stewart. An investigation into the effects of electric fields on Nephrops norvegicus. ICES Journal of Marine Science, 35(3):249–257, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/3/249/672749.

Steele:1975:OJB

[Ste75a] J. H. Steele. Obituaries: John B. Tait (7 June 1900–20 March 1973). ICES Journal of Marine Science, 36(2):99–100, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/36/2/99/746415.

Stewart:1975:CSE

[Ste75b] Peter A. M. Stewart. Catch selectivity by electrical fishing systems. ICES Journal of Marine Science, 36(2):106–109, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/36/
2/106/746263.

Stewart:1977:SRF

[Ste77] Peter A. M. Stewart. A study of the response of flatfish (Pleuronectidae) to electrical stimulation. ICES Journal of Marine Science, 37(2):123–129, February 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/2/123/603420.

Stirling:1972:PCE

[Sti72] Hadrian P. Stirling. The proximate composition of the European bass, Dicentrarchus labrax (L.) from the Bay of Naples. ICES Journal of Marine Science, 34(3):357–364, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/357/758469.

Studenetsky:1978:OGV

[Stu78] S. A. Studenetsky. Obituary: Georgi V. Nikolsky (6 May 1910– 2 February 1977). ICES Journal of Marine Science, 38(2): 129–130, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/129/743117.

Szekielda:1972:USS

[Sze72] Karl-Heinz Szekielda. Upwelling studies with satellites. ICES Journal of Marine Science, 34(3):379–388, October 1972. CO-DEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic).

URL http://academic.oup.com/icesjms/article/34/3/379/
758476.

Szekielda:1973:VOS

[Sze73] Karl-Heinz Szekielda. Validity of ocean surface temperatures monitored from satellites. ICES Journal of Marine Science, 35(1):78–86, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/78/769793.

Szekielda:1974:MAM

[Sze74] Karl-Heinz Szekielda. A multichannel approach to monitor sea surface temperature from space along the NW Coast of Africa. ICES Journal of Marine Science, 35(3):363–364, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/35/
3/363/672838.

Szekielda:1976:FTC

[Sze76] Karl-Heinz Szekielda. Fast temperature changes in the upwelling area along the NW coast of Africa. ICES Journal of Marine Science, 36(3):199–204, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.

com/icesjms/article/36/3/199/639213.

Tail:1974:R

[Tai74] J. B. Tail. Reviews. ICES Journal of Marine Science, 35(2): 216–217, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/216/696065.

Talbot:1977:DPE

[Tal77a] J. W. Talbot. The dispersal of plaice eggs and larvae in the Southern Bight of the North Sea. ICES Journal of Marine Science, 37 (3):221–248, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/221/637209.

Talbot:1977:R

[Tal77b] J. W. Talbot. Reviews. ICES Journal of Marine Science, 37(3): 319–320, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/319/637307.

Tangen:1976:DSS

[Tan76] Karl Tangen. A device for safe sedimentation using the Utermöhl technique. ICES Journal of Marine Science, 36(3):282–284, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
36/3/282/639261.

Talbot:1975:ARS

[TB75] J. W. Talbot and Myrtle J. Boon. The adsorption of rhodamine on to suspended sediments. ICES Journal of Marine Science, 36(2):128–132, May 1975. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/2/128/746287.

Throndsen:1970:NTW

[Thr70a] Jahn Throndsen. A non-toxic water sampler for shallow waters (0–50 m). ICES Journal of Marine Science, 33(2):298–300, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
33/2/298/800739.

Throndsen:1970:SSC

[Thr70b] Jahn Throndsen. A small sedimentation chamber for use on the wild M 40 and the Zeiss plankton microscopes. ICES Journal of Marine Science, 33(2):297–298, August 1970. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/33/2/297/800734.

Tiews:1976:PFM

[Tie76] Klaus Tiews. Paul Friedrich Meyer-Waarden. 5 December 1902– 15 February 1975. ICES Journal of Marine Science, 36(3): 197–198, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/197/639210.

Tambs-Lyche:1972:NTM

[TL72] Hans Tambs-Lyche. Notes on techniques and methods for seawater analysis. ICES Journal of Marine Science, 34(3):505, October 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
3/505/758553.

Treguer:1976:NTM

[TLC76] Paul Treguer, Pierre Le Corre, and Pierre Courtot. Notes on techniques and methods for sea water analysis: Dosage automatique des orthophosphates dans l’eau de mer utilisation de l’AutoAnalyzer(R) II. ICES Journal of Marine Science, 36(3): 289–294, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/289/639270.

Thompson:1974:MEG

[TM74] D. Thompson and S. Mostert. Muscle esterase genotypes in the pilchard, Sardinops ocellata. ICES Journal of Marine Science, 36 (1):50–53, November 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/1/50/642621.

Tomasini:1978:CAD

[TM78] J. A. Tomasini and J. Mazza. Comportement alimentaire de deux copépodes calanoides (Centropages typicus et Acartia clausi) en milieux nutritifs à une seule algue. (French) [Feeding behavior of two calanoid copepods (Centropages typicus and Acartia clausi) in nutrient media with a single alga]. ICES Journal of Marine Science, 38(2):154–179, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/154/743144.

Tomasini:1979:CAC

[TM79] J. A. Tomasini and J. Mazza. Comportement alimentaire de Centropages typicus (Copepoda, Calanoida) en milieux nutritifs á deux algues et en présence de nauplii d’Artemia sp.. (French) [Feeding behavior of Centropages typicus (Copepoda, Calanoida) in nutrient media á two algae and in the presence of Artemia sp]. ICES Journal of Marine Science, 38(3):293–307, August 1979. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://

academic.oup.com/icesjms/article/38/3/293/646590.

Ulltang:1977:SEL

[Ull77] Ø. Ulltang. Sources of errors in and limitations of Virtual Population Analysis (Cohort Analysis). ICES Journal of Marine Science, 37(3):249–260, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/249/637216.

VanBanning:1974:TRI

[Van74] P. Van Banning. Two remarkable infestations by Lernaeocera spp. (Copepoda parasitica). ICES Journal of Marine Science, 35 (2):205–206, February 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/35/2/205/696022.

vanBanning:1971:SNS

[vB71] P. van Banning. Some notes on a successful rearing of the herringworm, Anisakis marina L. (Nematoda: Heterocheilidae). ICES Journal of Marine Science, 34(1):84–88, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/34/1/84/602234.

Voipio:1978:IH

[Voi78] Aarno Voipio. Ilmo Hela. ICES Journal of Marine Science, 38(1):1– 2, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/1/630167.

vanOutryve:1976:IAC

[vOS76] E. van Outryve and P. Sorgeloos. An inexpensive automatic counter for cultures of aquatic organisms. ICES Journal of Marine Science, 36(3):284–286, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/284/639264.

Walker:1971:ESE

[Wal71] M. Greer Walker. Effect of starvation and exercise on the skeletal muscle fibres of the cod (Gadus morhua L.) and the coalfish (Gadus virens L.) respectively. ICES Journal of Marine Science, 33(3):421–427, May 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/33/3/421/711516.

Walne:1978:Ra

[Wal78a] P. R. Walne. Reviews. ICES Journal of Marine Science, 38(1):122, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
38/1/122/630189.

Walne:1978:Rb

[Wal78b] P. R. Walne. Reviews. ICES Journal of Marine Science, 38 (1):122–124, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/1/122/630192.

Watson:1977:UTC

[WBH77] A. Watson, C. Bye, and R. G. Houghton. The use of trawling in conjunction with plankton surveys as an aid to fish egg identification in the west central North Sea. ICES Journal of Marine Science, 37(3):310–313, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/37/3/310/637285.

Walker:1972:TNW

[WBP72] M. Greer Walker, A. C. Burd, and G. A. Pull. The total numbers of white skeletal muscle fibres in cross section as a character for stock separation in North Sea herring (Clupea harengus L.). ICES Journal of Marine Science, 34(2):238–243, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/238/595525.

Walne:1972:EPS

[WD72] P. R. Walne and G. J. Dean. Experiments on predation by the shore crab, Carcinus maenas L., on Mytilus and Mercenaria. ICES Journal of Marine Science, 34(2):190–199, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/190/595517.

Weihs:1978:TST

[Wei78] D. Weihs. Tidal stream transport as an efficient method for migration. ICES Journal of Marine Science, 38(1):92–99, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/38/
1/92/630297.

Wert:1977:R

[Wer77] Richard T. Wert. Reviews. ICES Journal of Marine Science, 37(3):320, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/37/3/320/637311.

Weston:1972:FSA

[Wes72] D. E. Weston. Fisheries significance of the acoustic attenuation due to fish. ICES Journal of Marine Science, 34(2):306–308, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
2/306/595539.

Wiborg:1976:LMM

[Wib76a] Kr. Fr. Wiborg. Larval mortality in marine fishes and the critical period concept. ICES Journal of Marine Science, 37(1):111, May 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
37/1/111/622161.

Wiborg:1976:FCE

[Wib76b] Kristian Fredrik Wiborg. Fishery and commercial exploitation of Calanus finmarchicus in Norway. ICES Journal of Marine Science, 36(3):251–258, February 1976. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/36/3/251/639243.

Wiebe:1972:FIR

[Wie72] Peter H. Wiebe. A field investigation of the relationship between length of tow, size of net and sampling error. ICES Journal of Marine Science, 34(2):268–275, March 1972. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.

oup.com/icesjms/article/34/2/268/595533.

Walker:1978:MPP

[WJA78] M. Greer Walker, F. R. Harden Jones, and G. P. Arnold. The movements of plaice (Pleuronectes platessa L.) tracked in the open sea. ICES Journal of Marine Science, 38(1):58–86, April 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL

http://academic.oup.com/icesjms/article/38/1/58/630257.

Walne:1978:CIO

[WM78] P. R. Walne and P. F. Millican. The condition index and organic content of small oyster spat. ICES Journal of Marine Science, 38 (2):230–233, November 1978. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (electronic). URL http://academic.oup.com/

icesjms/article/38/2/230/743218.

Wood:1971:PSA

[Woo71] E. J. Ferguson Wood. Phytoplankton study — an appraisal. ICES Journal of Marine Science, 34(1):123–126, October 1971. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289 (elec-

tronic). URL http://academic.oup.com/icesjms/article/34/
1/123/602157.

Walne:1974:EGF

[WS74] P. R. Walne and B. E. Spencer. Experiments on the growth and food conversion efficiency of the spat of Ostrea edulis L in a recirculation system. ICES Journal of Marine Science, 35(3):303–318, June 1974. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/3/303/672808.

Windom:1973:MNA

[WTS73] H. Windom, F. Taylor, and R. Stickney. Mercury in North Atlantic plankton. ICES Journal of Marine Science, 35(1):18–21, June 1973. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
35/1/18/769677.

Wootten:1977:SBL

[WW77] R. Wootten and I. F. Waddell. Studies on the biology of larval nematodes from the musculature of cod and whiting in Scottish waters. ICES Journal of Marine Science, 37(3):266–273, September 1977. CODEN ICESEC. ISSN 1054-3139 (print), 1095-9289

(electronic). URL http://academic.oup.com/icesjms/article/
37/3/266/637225.